### v1

# Seo Hyun Kim1     Kai Tzu-iunn Ong1 Taeyoon Kwon1     Namyoung Kim1     Keummin Ka1     SeongHyeon Bae1 Yohan Jo2     Seung-won Hwang2     Dongha Lee1     Jinyoung Yeo1

Abstract
Large language models (LLMs) are capable of processing lengthy dialogue histories during prolonged interaction with users without additional memory modules; however, their responses tend to overlook or incorrectly recall information from the past. In this paper, we revisit memory-augmented response generation in the era of LLMs. While prior work focuses on getting rid of outdated memories, we argue that such memories can provide contextual cues that help dialogue systems understand the development of past events and, therefore, benefit response generation. We present [Uncaptioned image]Theanine, a framework that augments LLMs’ response generation with memory timelines – series of memories that demonstrate the development and causality of relevant past events. Along with Theanine, we introduce TeaFarm, a counterfactual-driven question-answering pipeline addressing the limitation of G-Eval in long-term conversations. Supplementary videos of our methods and the [Uncaptioned image]TeaBag dataset for TeaFarm evaluation are in https://theanine-693b0.web.app/.
[Uncaptioned image]Theanine: Revisiting Memory Management in Long-term Conversations with Timeline-augmented Response Generation
Seo Hyun Kim1thanks: These two authors are the co-first authors and both have the right to list their names first on their CVs.     Kai Tzu-iunn Ong1 Taeyoon Kwon1thanks: T Kwon, N Kim, and K Ka are the co-second authors.     Namyoung Kim1     Keummin Ka1     SeongHyeon Bae1 Yohan Jo2     Seung-won Hwang2     Dongha Lee1     Jinyoung Yeo1 1Yonsei University, 2Seoul National University
1 Introduction
Recently, the impressive performance of large language models (LLMs) has caused a trend of having longer conversations with them across multiple sessions (Fui-Hoon Nah et al., 2023). Such prolonged interactions require the models to memorize previous events or speaker information, tailoring the responses accordingly to keep up user interest (Adiwardana et al., 2020). While the large context windows of LLMs [^1] Recent LLMs, such as GPT-4 (Achiam et al., 2023) and LLama3 (MetaAI, 2024), have contexts windows of 128,000 and 1,040,000 tokens, respectively. allow them to process the entire dialogue history for response generation (RG), this often leads to biased attention toward the latest user utterance, ignoring relevant contexts from the past (Liu et al., 2024), as presented in Figure 1 (a).
Figure 1: Empirical examples of failed responses due to (a) bias to the latest input (b) the absence of an important past event (“afraid of cruise ships”) on the timeline. (c) shows a response augmented with the memory timeline.
Figure 2: The overview of Refer to captionTheanine. Left: Linking new memories to the memory graph after finishing a dialogue session; Right: Memory timeline retrieval, refinement, and response generation in a new dialogue session.
One straightforward solution is to condense past conversations into summarized memories and retrieve them to augment response generation in later sessions (Xu et al., 2022a, b; Lu et al., 2023), but the growing span of memories can hinder the quality of memory retrieval as dialogue sessions accumulate. Although this can be mitigated by updating the older memories (Bae et al., 2022; Zhong et al., 2024; Kim et al., 2024a), we observe severe information loss. As illustrated in Figure 1 (b), an earlier memory on the timeline, which contains an important speaker persona (i.e., “afraid of ships”), is removed during memory update (thus can not be retrieved), yielding a suboptimal response.
Motivated by these, in this paper, we revisit memory-augmented response generation for long-term conversations and address the aforementioned bottlenecks. Specifically, we focus on using the timeline built with relevant memories to augment response generation. To this end, we present [Uncaptioned image]Theanine, [^2] L-theanine is an amino acid found in green tea that has been linked to memory improvement (Nguyen et al., 2019). a framework of timeline-augmented chain-of-thought reasoning for response generation in long-term conversations.
Inspired by how humans naturally link new memories to existing memories of relevant events and retrieve them as a whole (Bartlett, 1995), Theanine manages memories with a graph structure, i.e., the memory graph, where memories are linked based on how they relate to each other (Phase I). Intuitively, this can be done by linking each memory to others that have high text similarity. Yet, we find this naive approach not much beneficial to response generation, due to the plain connection between memories (e.g., “this happened $\rightarrow$ that occurred”). To resolve this, we leverage an LLM to dynamically link memories based on their temporal and cause-effect commonsense relations (Hwang et al., 2021) via our designed algorithm. Next, in memory retrieval for generating a new response (Phase II-1), instead of solely retrieving top-$k$ memories using current conversation as query, we retrieve the whole memory timeline representing the development of relevant events. Then, to tackle the discrepancy between off-line memory construction and online deployment (Phase II-2), Theanine refines the retrieved timeline such that it provides tailored information for the current conversation. Finally, in response generation (Phase III), we leverage LLMs’ chain-of-thought (CoT) reasoning ability (Wei et al., 2022) to reason over current conversation and refined timelines, conclude useful information, and generate the final response. Our contributions are three-fold:
  • We revisit memory-augmented response generation in long-term conversations with a novel timeline-augmented framework, [Uncaptioned image]Theanine. In both LLM-based (i.e., G-Eval) and human evaluation, Theanine demonstrates (i) better responses that are detailed/non-generic and (ii) efficiency in properly referencing past conversations.
  • The lack of ground-truth mapping between ongoing conversations (queries) and memories for retrieval poses a challenge in evaluating memory-augmented dialogue systems. We present TeaFarm, a counterfactual-driven question-answering pipeline, that assesses systems in referencing past conversations, which does not require human intervention.
  • We further curate [Uncaptioned image]TeaBag, the counterfactural question-answer data for TeaFarm evaluation on two widely used benchmark datasets of long-term conversations: Multi-Session Chat (MSC) (Xu et al., 2022a) and Conversation Chronicles (CC) (Jang et al., 2023).
2 Methodologies
We present [Uncaptioned image]Theanine (Figure 2), a framework that resembles the mechanism of how humans store and retrieve memories for conversations:
2.1 Memory Graph Construction (Phase I)
To replicate how humans construct memories, we approach memory management based on a graph structure, i.e., the memory graph $G$:
$G=(V,E)$ (1)
$V=\{m_{1},m_{2},...,m_{|V|}\}$ (2)
$m=(\textit{event},\textit{time})$ (3)
$E=\{\langle m_{i},r_{ij},m_{j}\rangle|m_{i},m_{j}\in V\land r_{ij}\in R\}$ (4)
$R=\{\texttt{Cause},\texttt{Reason},\texttt{Want},...,\texttt{SameTopic}\}$ (5)
In $G$, vertices $V$ are memories $m$ summarized from the conversations. Each memory $m=(\textit{event},\textit{time})$ consists of an event [^3] In this work, “event” denotes information perceived by the dialogue system, including things done/said by speakers and the acknowledgement of speaker personas. and the time it is formed (summarized). Each directed edge $e\in E$ between two connected $m$ indicates their temporal order and their cause-effect commonsense relation $r\in R$:
At the end of dialogue session $t$, [Uncaptioned image]Theanine starts linking each new memory $m_{new}$ summarized from session $t$ to the memory graph $G^{t}$.
Phase I-1: Identifying associative memories for memory linking.
Following how humans link new memories to existing ones that are related to a similar event/topic, i.e., the associative memories, Theanine starts by identifying these associative memories from the memory graph $G^{t}$.
Formally, given a newly-formed memory $m_{new}$ waiting to be stored, the associative memories $M_{a}$ of $m_{new}$ is defined as the set of $m_{i}\in G^{t}$ having top-$j$ text similarity with $m_{new}$ (i.e., $|M_{a}|=j$).
Phase I-2: Relation-aware memory linking.
Intuitively, we can link $m_{new}$ to $m\in M_{a}$ using edges that indicate their text similarity and chronological order, we find such simplified connection (e.g., “this happened $\rightarrow$ that similar event occurred”) can yield a context-poor graph that does not help response generation much (Section 4).
Humans, on the other hand, interpret events by considering the relation between them, such as “how does an event affect the other?” or “why did this person make that change?”.
Therefore, we adopt a relation-aware memory linking, where each edge between two memories is encoded with their cause-effect commonsense relation $r\in R$, along with their temporal order. These relations are adopted from Hwang et al. (2021) and modified by us, including HinderedBy, Cause, Want, and 4 more (Appendix A.1).
We start by determining the relation between $m_{new}$ and each associative memory. Formally, for each pair of $m_{new}$ and $m\in M_{a}$, the LLM assigns a relation $r\in R$ based on their event, time and their origin conversations:
$M_{a}^{*}=\{m_{i}\in M_{a}\mid\Upsilon(m_{i},m_{new})\in R\}$ (6)
where $\Upsilon(\cdot,m_{new})\in R$ indicates that the given memory is assigned with an $r\in R$ with $m_{new}$, [^4] Limited by the performance of retrievers, it is possible that an $m\in M_{a}$ does not have a relation, other than just textual overlap, with $m_{new}$. We address this by allowing the LLM to output “None”. and such assigned memories are defined as $M_{a}^{*}$.
Figure 3: Locating memories to be linked to $m_{new}$.
We then proceed to link $m_{new}$ to the graph, a straightforward approach is linking $m_{new}$ to all $m\in M_{a}^{*}$. Yet, we find it algorithmically inefficient (Figure 7). Thus, we selectively link $m_{new}$ to only those $m\in M_{a}^{*}$ that are recent. To identify those recent memories, we first locate every connected component $C_{i}\subset G^{t}$ that contains at least one $m\in M_{a}^{*}$, as shown in Figure 3 (a) and (b):
=$\mathbb{C}=\{C_{i}\subset G^{t}\mid\mathtt{V}(C_{i})\cap M_{a}^{*}\neq\emptyset\ \}$ (7)
where $\mathbb{C}$ is the collection of those $C$ and $\mathtt{V}(\cdot)$ represents “vertices in”. Now, we link $m_{new}$ to the most recent $m\in M_{a}^{*}$ in each $C_{i}\subset\mathbb{C}$ (Figure 3 (c)). Therefore, the memories $M_{linked}$ that are linked to $m_{new}$ is defined as follows:
$M_{linked}=\{\Omega(\mathtt{V}(C_{i})\cap M_{a}^{*})\mid C_{i}\subset\mathbb{C}\}$ (8)
where $\Omega(\cdot)$ indicates “the most recent memory in”. Linking all memories from session $t$ to $G^{t}$, we then obtain a new memory graph $G^{t+1}$. The pseudo algorithm for Phase I is in Algorithm 1.
2.2 Timeline Retrieval and Timeline Refinement (Phase II)
Drawing motivation from Figure 1, [Uncaptioned image]Theanine augments response generation (RG) with timelines of relevant events to address the information loss caused by conventional memory management (Xu et al., 2022a; Bae et al., 2022). With $G^{t+1}$, Theanine performs these steps for RG in session $t+1$:
Preparation: Top-k memory retrieval.
During the conversation, using the current dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ of $n$ utterances $u$ as query, we first retrieve top-$k$ memories $M_{re}=\{m_{re1},...,m_{rek}\}$.
Phase II-1: Retrieving raw memory timelines.
After that, we wish to also access memories that are centered around relevant events. Formally, given an $m_{re}\in M_{re}$, we further retrieve the connected component $C_{re}\subset G^{t+1}$ that contains $m_{re}$.
Since this collection of memories (i.e., $C_{re}$) can be “tangled up” together (i.e., connected in a complex manner) due to the graph structure, we proceed to untangle it into several memory timelines, each representing a series of events about $m_{re}$ that starts out similarly yet branches into slightly different development. For that, we first locate the earliest memory in $C_{re}$ as a starting point $m_{start}$ for all timelines, as shown in Figure 4 (left).
$m_{start}=\Theta(\mathtt{V}(C_{re}))$ (9)
where $\Theta$ indicates “the oldest memory in”
Figure 4: Extracting raw memory timelines $\tau$ from the retrieved connected component $C_{re}$.
Next, starting from $m_{start}$, we untangle the memories by tracing through the future direction and extract every possible linear graph containing $m_{re}$ (two in Figure 4) from $C_{re}$, until reaching an endpoint $\tau[-1]$ with an out-degree of 0 (i.e., $deg^{+}(\tau[-1])=0$, which means no directed edge goes out from it). Each of them is considered a raw memory timeline $\tau$, demonstrating a version of the evolution of $m_{re}$ and its relevant events:
$\begin{split}\mathcal{T}={}&\{\tau\subset C_{re}\mid\tau\smash{{\text{ is a directed linear}}}\\ &\smash{{\text{ graph s.t. }}}m_{start},m_{re}\in\tau\ \\ &\land deg^{+}(\tau[-1])=0\}\end{split}$ (10)
We then sample $n$ raw timelines $\tau$ from $\mathcal{T}$ [^5] Based on our experiments, we empirically set $n$ to 1. Repeating Phase II-1 for all retrieved top-$k$ memories, we collect a set of retrieved raw memory timelines $\mathbb{T}=\cup\,\mathcal{T}$, where $|\mathbb{T}|=k^{*}n$.
Phase II-2: Context-aware timeline refinement.
Although we have constructed the memory graph using temporal and commonsense relations to improve its informativeness, directly applying the retrieved timelines for response generation can be suboptimal (RQ3, Section 4), because graph construction does not take the current conversation into consideration, i.e., they are constructed off-line.
In this phase, [Uncaptioned image]Theanine tackles such a discrepancy between off-line memory construction and online deployment (i.e., an ongoing current conversation) via a context-aware timeline refinement. Motivated how LLMs can self-refine their previous generation (Madaan et al., 2024). We leverage LLMs to refine raw timelines into a rich resource of information crafted for the current conversation, by removing redundant information or highlighting information that can come in handy. Formally, given the current dialogue $\mathcal{D}$ and the retrieved raw memory timelines $\mathbb{T}$, the LLM tailors all $\tau\in\mathbb{T}$ into refined timelines $\mathbb{T}_{\Phi}$:
$\mathbb{T}_{\Phi}=\{\underset{\tau_{\Phi}}{\text{argmax}}\,P_{\text{LLM}}(\tau_{\Phi}|\mathcal{D},\tau)\mid\tau\in\mathbb{T}\}$ (11)
All refined timelines $\mathbb{T}_{\Phi}$ are then used to augment the response generation. We provide the pseudo algorithm for Phase II in Algorithm 2.
2.3 Timeline-augmented Response Generation (Phase III)
Finally, we move on to response generation. Here, [Uncaptioned image]Theanine leverages LLMs’ chain-of-thought (CoT) reasoning to discover user intention from the current conversation (Chae et al., 2023) and selectively extract useful information provided in the refined timelines. Formally, given the current dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ and refined timelines $\mathbb{T}_{\Phi}$, the LLM sequentially generates a CoT rationale $\mathcal{R}$ demonstrating how it comprehends $\mathcal{D}$ and $\mathbb{T}_{\Phi}$, and the next response $\bar{u}_{t+1}$:
¯$\bar{\mathcal{R}}=\underset{\mathcal{R}}{\text{argmax}}\,P_{\text{LLM}}(\mathcal{R}|\mathcal{D},\mathbb{T}_{\Phi})$ (12)
$\Rightarrow\bar{u}_{n+1}=\underset{u_{n+1}}{\text{argmax}}\,P_{\text{LLM}}({u}_{n+1}|\mathcal{D},\mathbb{T}_{\Phi},\bar{\mathcal{R}})$ (13)
where $\Rightarrow$ denotes sequential text generation.
3 Experimental Setups
3.1 Datasets of Long-term Conversations
There are few datasets of long-term conversations. First, Multi-Session Chat (MSC) (Xu et al., 2022a), which is built upon the conversations from Persona-Chat (Zhang et al., 2018) by extending their follow-up conversations throughout multiple sessions. Soon after MSC, DuLeMon (Xu et al., 2022b) and CareCall (Bae et al., 2022) are proposed for long-term conversations in Mandarin and Korean, respectively. Recently, Jang et al. (2023) release a new multi-session datasets, Conversation Chronicles (CC). Different from MSC, speakers in CC are augmented with diverse relationships, such as “employee and boss”. Apart from the above open-domain datasets, there is also Psychological QA, [^6] https://www.xinli001.com/ which addresses long-term conversation in Mandarin regarding clinical scenarios.
In this work, we adopt MSC and CC for evaluation. Since we focus on English conversations, we do not include DuleMon, CareCall, and Psychological QA. We leave such multilingual settings and clinical scenarios to future work.
Datasets: Multi-session Chat (MSC) & Conversation Chronicles (CC)
Dialogue Session: Session 3 Session 4 Session 5 Avg.
[Uncaptioned image]Theanine Theanine’s Winning Percentage
          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
All Dialogue History 71% 82% 77% 73% 81% 77% 72% 79% 77% 72% 81% 77%
All Memories and $\mathcal{D}$ 69% 64% 66% 69% 69% 68% 70% 68% 68% 69% 67% 67%
     + Memory Update 56% 57% 60% 56% 56% 58% 56% 58% 60% 56% 57% 60%
Memory Retrieval 67% 62% 68% 72% 67% 68% 72% 65% 68% 70% 65% 68%
     + Memory Update 57% 56% 59% 58% 57% 61% 55% 58% 59% 57% 57% 60%
RSum-LLM 38% 61% 69% 35% 57% 67% 37% 57% 65% 37% 58% 67%
MemoChat 60% 74% 71% 60% 69% 72% 60% 72% 71% 60% 72% 71%
COMEDY 66% 54% 78% 64% 54% 77% 62% 53% 76% 64% 53% 77%
Golden Responses 63% 76% 61% 63% 73% 61% 62% 72% 61% 63% 73% 61%
Table 1: Win rates of Theanine in head-to-head G-Eval (of all generated responses) with baselines. Spe = specificity; Mem = memorability; Cst = consistency. Color green denotes metrics regarding the reference of memories. Color gray shows where ours loses more. Underline: p-value $<0.001$. *: $0.001<$ p-value $<0.05$.
3.2 Baselines
To evaluate Theanine in long-term conversation, besides naive settings using all past dialogues or memories, we include the following baselines.
Memory retrieval. Following Xu et al. (2022a) and Xu et al. (2022b), this setting uses a retriever to retrieve memories with the current dialogue context to augment response generation.
+ Memory update. We use LLMs to perform the memory updating mechanism developed by Bae et al. (2022) to update the memory pool at the end of each session to enhance the above baselines.
RSum-LLM. Proposed by Wang et al. (2023), it uses an LLM to (i) recursively summarize and update the memory pool at the end of each session, (ii) select top memories based on the current dialogue, and (iii) generate responses.
MemoChat. Proposed by Lu et al. (2023), it uses LLMs’ CoT reasoning ability to (i) conclude important memories from past conversations in a structured topic-summary-dialogue manner, (ii) select memories, and (ii) generate responses.
COMEDY. Proposed by Chen et al. (2024b), rather than memory retrieval, it first summarizes session-level past summaries and then compresses all of them into short sentences of events, user portraits (behavior patterns, emotion, etc.) and user-bot relation, using them to augment response generation.
3.3 Models and Implementation Details
Large language models. In all experiments including baselines, we adopt ChatGPT (gpt-3.5-turbo-0125) from OpenAI (2023) for (i) memory summarization, (ii) memory update, and (iii) response generation. We set the temperature to $0.75$.
Retrievers. We use text-embedding-3-small (OpenAI, 2024b) from OpenAI to calculate the similarity of texts. For the identification of top-$j$ associative memories (Phase I-1) and top-$k$ memory retrieval (Phase II Preparation), we set $j$ and $k$ to 3. For the “Memory Retrieval” baseline, we retrieve 6 memories following Xu et al. (2022a).
Dialogue sessions. We perform evaluation using sessions 3-5 from both datasets, as all settings are almost identical before the end of session 2 (there is no memory to update before then).
4 Evaluation Scheme 1: Machine and Human Evaluation
To evaluate [Uncaptioned image]Theanine, we first start with G-Eval (Liu et al., 2023), an LLM-based framework for assessing LMs’ responses, and human evaluation. [^7] Details for G-Eval and human evaluation are in Appendix D and F. Dataset-specific results are in Appendix E. Here, we focus on the following research questions (RQs):
RQ1. Can Theanine generate context-rich responses, rather than those generic and plain?
RQ2. Can Theanine reference past conversations better than other methods?
RQ3. Do phases in Theanine make sense?
RQ4. Do humans agree with Theanine’s procedures, and how do they think of the final responses?
[Uncaptioned image]Theanine yields responses that are more detail-rich (RQ1).
In Table 1, we first assess whether generated responses are detail-rich, rather than generic and plain, i.e., specificity (Spe). In both datasets, Theanine yields significantly high win rates in the head-to-head comparison with all baselines except for RSum-LLM. We presume it is due to the fact that RSum-LLM tends to generate much longer responses (avg. 77 words per response) than other settings (ours = 32 words; avg of other baselines = 30 words).
[Uncaptioned image]Theanine utilizes past memories efficiently and correctly (RQ2).
In long-term conversations, it is crucial and necessary to further assess (i) whether the dialogue system utilizes information from the past when necessary (i.e., memorability) and make sure (ii) the response does not contradict previous conversations (i.e., consistency). We present the results in Table 1. Theanine demonstrates high win rates in the comparisons with baselines. However, COMEDY performs almost as well as us regarding memorability, which we assume is because it always applies two additional LLMs to extract and compress past conversations into concise memos before each response generation (no retrieval). However, as we expected, always referencing concise and recent user information can lead to contradiction to older information, thus clearly beaten by Theanine in terms of consistency (Cst).
Ablation results: Phases in [Uncaptioned image]Theanine matter (RQ3).
Table 2 presents the results where key phases in Theanine are removed. We observe that the removal of relation-aware memory linking (Phase I-2) and context-aware timeline refinement (Phase II-2) leads to large performance drops regarding the reference of past information (i.e., the original Theanine shows higher win rates), especially when timeline refinement is absent.
This justifies our methodological designs: (i) In graph construction, linking memories with cause-effect relation benefits RG more than naive connection using only text similarity and chronological order; (ii) By further refining memories built off-line (i.e., without considering current conversations) before RG, we can elicit contextual cues that are tailored for the current conversation.
Multi-session Chat (Session 3-5)
[Uncaptioned image]Theanine Ours’ Win Rate
          vs. Mem Cst Avg.
w/o Relation-aware Linking 62% 65% 64%
w/o Timeline Refinement 62% 69% 65%
Conversation Chronicles (Session 3-5)
w/o Relation-aware Linking 60% 65% 63%
w/o Timeline Refinement 64% 68% 66%
Table 2: Ablation results (p-values $<0.001$).
Figure 5: Results of human evaluation.
Humans think [Uncaptioned image]Theanine makes sense and properly references past conversations.
As reported in Figure 5, human judges largely agree (92%) that Theanine’s relation-aware memory linking is performed appropriately, and they all agree (100%) that timeline refinement can elicit more helpful/tailored information for response generation. Most importantly, Theanine produces responses that either entail (i.e., faithfully reflect) past conversations (34%) or, at worse, neutral statements (57%) in 91% of the cases from human standpoints, manifesting its’ efficacy in referencing past information in long-term conversations.
Examples of [Uncaptioned image]Theanine in terms of each phase are in Appendix G.
5 Evaluation Scheme 2: TeaFarm – a Counterfactual-driven Evaluation Pipeline for Long-term Conversations
Despite the success of G-Eval (Liu et al., 2023) in evaluating responses of LMs, accessing dialogue systems’ ability to utilize past memories in long-term conversation still remains challenging. This is due to the unavailability of ground-truth mapping between current conversations and the correct memories for retrieval. Although we can resolve this by feeding the evaluator LLM (e.g., GPT-4) the entire memory pool or past history and prompt it to determine whether a response correctly recalls past conversations, the evaluation can be limited by the performance of the evaluator LLM itself.
To overcome this, along with Theanine, we present TeaFarm, a counterfactual-driven pipeline for evaluating memory-augmented response generation in long-term conversations.
5.1 Testing Dialogue Systems’ Memory via Counterfactual Questions
In TeaFarm, we proceed to “trick” the dialogue systems into generating incorrect responses, and the system must correctly reference past conversations to avoid being misled by us. Specifically, we talk to the dialogue system while acting as if a non-factual statement is true (thus counterfactual). Figure 6 presents some examples of counterfactual questions and the corresponding facts.
Figure 6: Examples of counterfactual questions.
In practice, when we want to evaluate a system that has been interacting with the user for multiple sessions, we first (1) collect all past conversations and summarize them session by session. Then, we (2) feed the LLM [^8] We apply GPT-4 (gpt-4) with a temperature of $0.75$. the collected summaries in chronological order such that it can capture the current stage of each discussed event, e.g., “Speaker B has never been to Japan at this moment”, and (3) generate counterfactual questions from the perspective of both speakers (and the correct answers). After that, we (4) kick off a new conversation, chat for a while, then (5) naturally ask the counterfactual question, and (6) assess the correctness of its response. Overview illustration and prompts for TeaFarm are in Appendix B and H.
5.2 TeaFarm Results
Settings / Datasets MSC CC Avg.
Memory Retrieval 0.16 0.19 0.18
+ Memory Update 0.16 0.19 0.18
RSum-LLM 0.04 0.08 0.06
MemoChat 0.09 0.15 0.12
COMEDY 0.06 0.18 0.12
[Uncaptioned image]Theanine 0.17 0.24 0.21
      w/o Relation-aware Linking 0.06 0.13 0.10
      w/o Timeline Refinement 0.15 0.16 0.16
Table 3: TeaFarm results regarding correct/total percentages of model responses. Tested in conversations from MSC and CC with 200 counterfactual questions.
In Table 3, we present model performance in TeaFarm evaluations. Theanine outperforms all baselines, especially in CC. Ablations perform worse than original Theanine, again proving the efficacy and necessity of relation-aware linking and timeline refinement. Surprisingly, all settings result in low correct rates, qualifying TeaFarm as a proper challenge for stress-testing dialogue systems in long-term conversations.
Interestingly, baselines that utilize retrievers (just like Theanine) show superior performance than those relying on LLMs for memory selection (i.e., RSum-LLM, MemoChat, and COMEDY). [^9] Memory update does not affect Memory Retrieval’s performance. We believe it is because the counterfactual questions are made to counter the newest stage of each event. The removal of older memories thus does not have much impact. This pattern, unexpectedly, supports our idea of revisiting memory retrieval in the era of LLMs.
5.3 Dataset for TeaFarm: [Uncaptioned image]TeaBag
As a byproduct of TeaFarm, we curate [Uncaptioned image]TeaBag, a dataset for TeaFarm evaluation on MSC and CC. TeaBag consists of (1) 100 episodes of original conversations from MSC and CC (session 1-5), (2) follow-up conversations (session 6), where the speaker will naturally converse toward the counterfactual questions, and (3) the corresponding answers. TeaBag serve as a useful tool for stress-testing if a system can correctly reference past conversations. Dataset details are in Appendix C.
6 Further Discussions
Cost efficiency.
In Phase I-2, we choose to link $m_{new}$ to only the most recent associative memories in each $C_{i}$, rather than all associative memories in $M_{a}$. Figure 7 compares Theanine with Theanine-All, where $m_{new}$ is linked to all $m\in M_{a}$. We find that although Theanine-All costs 25% more, it is beaten by original Theanine. We believe this is because linking $m_{new}$ to all $M_{a}$ leads to more short timelines (e.g., an $m_{new}$ from session 4 is directly linked to an old memory from session 1), which provide less information than timelines built with relevant memories from more sessions. Regardless, it still yields lower loss rates than strong baselines. An open question here is whether we can reach a balance between cost and performance. We presume we can further improve the cost efficiency of Theanine in the long-term by training a memory linker using the collected memory-relation pairs from LLMs via knowledge distillation. We leave it to future work.
Growing span of memories.
Another inquiry left might be whether the growing span of memory would hinder memory retrieval in Theanine if there ever were hundreds of sessions. Although this can be a serious issue for conventional memory retrieval, we argue that it will be much mitigated in Theanine because: (i) We retrieve relevant memories as a whole in the form of timelines. This works as a safety net in scenarios where an important memory is missed out in top-$k$ retrieval – it might still be retrieved along with the timeline; (ii) We perform timeline refinements. Similarly, this also acts as a second insurance against mis-retrieved memories that are actually not helpful.
Figure 7: API cost and G-Eval results. Red shows baselines that lose even more than Theanine-All in the head-to-head comparisons with original Refer to captionTheanine.
7 Related Work
Long-term Conversations.
Since the release of MSC (Xu et al., 2022a), there have been many studies on long-term conversations. For instance, Bae et al. (2022) train a classifier as the memory updater to handle outdated memories in phone call scenarios. Recently, based on LLMs’ outstanding in-context learning ability, Wang et al. (2023) leverage LLMs to write, update, and select memories for response generation. Apart from LLMs’ power, human behaviors have also fostered methods for long-term conversations. For example, Zhong et al. (2024) apply humans’ forgetting curve to make memories that have been discussed stay in the memory pool for a longer time. Also drawing inspiration from how humans manage memories, our [Uncaptioned image]Theanine, on the other hand, addresses the connection between relevant events in both memory construction and retrieval, improving response quality in long-term conversations, which has never been explored to the best of our knowledge.
Memory-augmentation for personalized dialogue systems.
The trend of long-term interaction with dialogue systems promotes the adaptation of them for personalized needs (Chen et al., 2024a, c). As a pioneer, Xu et al. (2022b) train a persona extractor to create user-based memories in long-term conversations. However, training a personalized system for long-term use can be non-trivial due to the lack of data (Tseng et al., 2024). As a solution, Kim et al. (2024a) apply commonsense models and LLMs to augment existing long-term conversation datasets with high-quality persona sentences; Chen et al. (2024b) present a training-free LLM-based framework that extracts user behaviors from past conversations for personalized responses. Upon the success of LLMs, [Uncaptioned image]Theanine leverages them to build memory timelines, which represent the development of user personas/behaviors throughout the interaction, providing rich cues for personalized responses.
8 Conclusions
This work revisits memory management in multi-session/long-term conversations in the era of LLMs. Following how humans store and retrieve relevant memories as a whole, we present [Uncaptioned image]Theanine, a framework of timeline-augmented response generation, which elicits responses that are detail-rich and, more importantly, able to correctly reference previous conversations when necessary. We also propose a novel evaluation pipeline (TeaFarm) that addresses the limitation of G-Eval in long-term conversations as well as the dataset for it ([Uncaptioned image]TeaBag). We expect this paper to serve as a groundwork and inspiration for exploring more paradigms of memory construction/retrieval in dialogue systems.
9 Limitations
This work has the following limitations: First, the maximum amount of dialogue sessions in this study is limited to five due to the lack of longer open-domain English datasets of multi-session conversations. Hence, Theanine’s performance in longer conversations has not been investigated. As we mentioned in Section 6, we presume that its’ effectiveness can still hold true to some degree even in such scenarios. Yet, we do acknowledge the need to apply additional modules that directly address the growing span of dialogue history/memories. One possible way is to introduce the summarize-then-compress paradigm in COMEDY (Chen et al., 2024b) to Theanine, which further compresses session-level summaries into a combined short user/event description.
Second, retrieving and refining timelines for response generation can be less computationally efficient than conventional memory retrieval, as there might exist scenarios where timelines are not necessary, e.g., when there is no need to reference past conversations. We believe this can be mitigated by incorporating an additional selection logic/module to choose between memory retrieval, timeline retrieval, and no retrieval at all.
Lastly, although we include many recent frameworks as baselines, we failed to compare Theanine with MemoryBank (Zhong et al., 2024), a framework inspired by Ebbinghaus’s forgetting curve. This is because the time intervals between sessions in MSC and CC are either mostly measured in hours or not clearly specified (e.g., “a few months later”), whereas MemoryBank requires precise time intervals in days to apply the forgetting curve. Also, data used for MemoryBank focuses on Chinese clinical scenarios, making it not feasible for our study. However, we remain positive about applying such a mechanism to improve Theanine in our ongoing research.
10 Ethical Statements
LLMs might generate harmful, biased, offensive, sexual content. Authors avoid such content from appearing in this paper. We guarantee fair compensation for human evaluators from Amazon Mechanical Turk. We ensure an effective pay rate higher than 20\$ per hour based on the estimated time required to complete the tasks.
11 Acknowledgments
This work was partly supported by an IITP grant funded by the Korean Government (MSIT) (No. RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)). Jinyoung Yeo is the corresponding author (jinyeo@yonsei.ac.kr).
Appendix A Further Implementation Details
A.1 Cause-effect Commonsense Relations
We adopt and modify commonsense relations from Hwang et al. (2021) for our relation-aware memory linking. Below is the list of our commonsense relations $R$:
Changed: Events in A changed to events in B.
Cause: Events in A caused events in B.
Reason: Events in A are due to events in B.
HinderedBy: When events in B can be hindered by events in A, and vice versa.
React: When, as a result of events in A, the subject feels as mentioned in B.
Want: When, as a result of events in A, the subject wants events in B to happen.
SameTopic: When the specific topic addressed in A is also discussed in B.
Limited by the performance of retrievers, it is possible that an $m\in M_{a}$ does not have a relation, other than just textual overlap, with $m_{new}$. We address this by allowing the LLM to output None.
A.2 Algorithms for [Uncaptioned image]Theanine
The pseudo algorithms for Phase I and II are provided in Algorithm 1 and 2.
A.3 Computational Experiments
All computational experiments in this work are based on OpenAI API (OpenAI, 2024a). Thus, no computing infrastructure is required in this work.
Appendix B TeaFarm Evaluation
The overview of TeaFarm is illustrated in Figure 8.
Appendix C The [Uncaptioned image]TeaBag Dataset
We curate [Uncaptioned image]TeaBag, a dataset for TeaFarm evaluation on MSC and CC. TeaBag consists of:
  • 100 episodes of original conversations from Multi-Session Chat and Conversational Chronicles (session 1-5; 50 episodes from each dataset)
  • Two pairs of counterfactual QAs for each episode (200 pairs in total).
  • Two synthesized session 6 for each episode (thus 200 in total), each of which naturally leads to one of the counterfactual questions.
This dataset is made with GPT-4. The prompt for generation is in Appendix H. We expect future work to apply TeaBag to stress-test if their dialogue system can correctly reference past conversations.
TeaBag does not contain personally identifying information, as it is generated based on datasets where all contents are pure artificial creation, rather than contents collected from the real-world. Also, we have tried our best to confirm that this dataset does not contain any offensive content.
Appendix D G-Eval
G-Eval (Liu et al., 2023) is a framework using LLMs with chain-of-thoughts (CoT) and a form-filling paradigm, to assess the quality of models’ text generation. G-Eval with GPT-4 has been shown to generate evaluation results that highly align with human judgement (Liu et al., 2023; Kim et al., 2024b). Due to the high API cost for it, we conduct head-to-head comparisons between Theanine and 8 baselines using 100 episodes from MSC and CC (around 3,600 Theanine’s responses being compared in each comparison with one baseline), rather than the whole test sets. The prompts for G-Eval are in Figure 2021, and 22. We use SciPy to calculate p-values. [^10] https://scipy.org/
Appendix E Dataset-Specific G-Eval Results
We provide dataset-specific results of G-Eval in Table 4 and 5.
Appendix F Human Evaluation
We conduct human evaluation, with workers from Amazon Mechanical Turk (AMT). We construct the following three evaluations:
  • Appropriateness of relation-aware memory linking: In this evaluation, we ask the workers to judge whether they agree that the relation-aware linking is properly done for two given memories.
  • Helpfulness of context-aware timeline refinement: This evaluation requires the workers to determine if they agree that our context-aware refinement really tailors a raw timeline into a resource of useful information for generating the next response.
  • The quality of Theanine’s responses: Here, the workers are asked to judge if the responses of Theanine correctly refer to past conversations. After reading our responses and past memories, they choose whether the responses entail, contradict, or are neutral to past memories. To improve evaluation quality, we use GPT-4 to select responses for this specific evaluation based on past memories, addressing the fact that not every turn in the conversation requires previous information to generate the next response (In the other two evaluations, the samples are randomly selected).
Each data sample is judged by 3 different workers, and we report the results based on the majority rule. In the third evaluation, when every option (entailment, neutral, contradiction) gets one vote, we consider it neutral (13 samples in total). These human evaluations are conducted on 300 samples. The interfaces provided to AMT workers, which include detailed instructions for human evaluation, are shown in Figure 910, and 11.
Appendix G Empirical Examples
We provide several empirical examples of [Uncaptioned image]Theanine. Examples of relation-aware memory linking are in Figure 1213, and 14. Examples of utilizing refined timeline for response generation are in Figure 15.
Appendix H Prompts
The following are all prompts utilized in our study:
  • Relation-aware memory linking (Phase I-2): Figure 16.
  • Context-aware timeline refinement (Phase II-2): Figure 17.
  • Timeline-augmented Response generation (Phase III): Figure 18.
  • + Memory update (baseline): Figure 19.
  • RSum-LLM (baseline): We adopt the original prompt from Wang et al. (2023).
  • MemoChat (baseline): We adopt the original prompt from Lu et al. (2023).
  • COMEDY (baseline): We adopt the original prompt from Chen et al. (2024b).
  • G-Eval: Figure 20,  21, and  22.
  • Generating counterfactual QA in TeaFarm: Figure 23.
  • Generating session 6 in TeaFarm: Figure 24.
  • Evaluating model responses in TeaFarm: Figure 25.
Appendix I Terms for Use of Artifacts
We adopt the MSC and CC datasets from Xu et al. (2022a) and Jang et al. (2023), respectively. Both of these datasets are open-sourced for academic and non-commercial use. Our curated dataset, TeaBag, which will be released after acceptance, is open to academic and non-commercial use.
Algorithm 1 Memory Graph Construction (Phase I)
Require: Memory graph $G^{t}=(V^{t},E^{t})$
Require: New memories $M_{new}=\{m_{new1},...,m_{newN}\}$
Require: Set of relations $R=\{\texttt{Cause},\texttt{Reason},\texttt{Want},...,\texttt{SameTopic}\}$
Ensure: Memory graph $G^{t+1}=(V^{t+1},E^{t+1})$
1:$\Upsilon(m_{i},m_{j})=\begin{cases}r_{i,j},\text{if }m_{i}\text{ is assigned with }r_{i,j}\in R\text{ with }m_{j}\\ \texttt{None},\text{otherwise}\end{cases}$
2:$\Omega(V)=(\text{the most recent memory m}\in V)$
3:$E_{t+1}\leftarrow E_{t}$
4:for $m_{new}\in M_{new}$ do
5:     $M_{a}\leftarrow\{m_{i}\in V^{t}\mid m_{i}\text{ has top-{j} similarity with }m_{new}\}$
6:     $M_{a}^{*}\leftarrow\{m_{i}\in M_{a}\mid\Upsilon(m_{i},m_{new})=r\text{ for }r\in R\}$
7:     $\mathbb{C}\leftarrow\{C_{i}\mid C_{i}\text{ connected component of }G^{t}\text{ s.t. }\mathtt{V}(C_{i})\cap M_{a}^{*}\neq\emptyset\ \}$
8:     $M_{linked}\leftarrow\{\Omega(\mathtt{V}(C_{i})\cap M_{a}^{*})\mid C_{i}\in\mathbb{C}\}$
9:     $E_{new}\leftarrow\{\langle m_{i},\Upsilon(m_{i},m_{new}),m_{new}\rangle\mid m_{i}\in M_{linked}\}$
10:     $E_{t+1}\leftarrow E_{t+1}+E_{new}$
11:end for
12:$V^{t+1}\leftarrow V^{t}+M_{new}$
13:$G^{t+1}\leftarrow(V^{t+1},E^{t+1})$
14:return $G^{t+1}$
Algorithm 2 Timeline Retrieval and Timeline Refinement (Phase II)
Require: Memory graph $G=(V,E)$
Require: Dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$
Ensure: Collection of refined timelines $\mathbb{T}_{\Phi}$
1:$\Theta(V)=(\text{the oldest memory m}\in V)$
2:$M_{re}\leftarrow\{m_{i}\in V\mid m_{i}\text{ has top-{k} similarity with }D\}$
3:$\mathbb{C}_{re}\leftarrow\{C_{re}\mid C_{re}\text{ connected component of }G\text{ s.t. }\mathtt{V}(C_{re})\cap M_{re}\neq\emptyset\}$
4:$\mathbb{T}\leftarrow\{\}$
5:for $C_{re}\in\mathbb{C}_{re}$ do
6:     $m_{start}\leftarrow\Theta(\mathtt{V}(C_{re}))$
7:     $\mathcal{T}=\{\tau\subset C_{re}\mid\tau\text{ is a directed linear graph s.t.}m_{start},m_{re}\in\tau\ \land deg^{+}(\tau[-1])=0\}$
8:     $\mathbb{T}\leftarrow\mathbb{T}+\text{RandomSelection}(\mathcal{T})$
9:end for
10:$\mathbb{T}_{\Phi}\leftarrow\{\underset{\mathcal{T}_{\Phi}}{\text{argmax}}\,P_{\text{LLM}}(\mathcal{T}_{\Phi}|\mathcal{D},\tau)\mid\tau\in\mathbb{T}\}$
11:return $\mathbb{T}_{\Phi}$
Figure 8: The overview of TeaFarm Evaluation.
Datasets: Multi-session Chat (MSC)
Dialogue Session: Session 3 Session 4 Session 5 Avg.
[Uncaptioned image]Theanine Theanine’s Winning Percentage
          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
All Dialogue History 74% 79% 72% 78% 80% 74% 74% 76% 71% 75% 78% 72%
All Memories and $\mathcal{D}$ 71% 66% 64% 72% 69% 67% 70% 65% 63% 72% 66% 65%
     + Memory Update 59% 54%* 58% 58% 53% 58% 56%* 54%* 57%* 58% 54%* 58%
Memory Retrieval 68% 61% 65% 73% 67% 67% 72% 64% 65% 71% 64% 66%
     + Memory Update 56%* 53% 57% 61% 58% 61% 53%* 53% 53% 57% 55% 57%
RSum-LLM 44% 57% 68% 41% 54%* 67% 43% 53% 63% 42% 55% 66%
MemoChat 65% 75% 70% 64% 69% 72% 65% 74% 72% 65% 73% 71%
COMEDY 72% 64% 84% 72% 61% 84% 69% 60% 81% 71% 61% 83%
Table 4: Winning percentage of Theanine in head-to-head G-Eval (of all generated responses) with baselines. Tested with Multi-Session Chat (MSC). Color gray shows where ours loses more. Underline: p-value $<0.001$. *: $0.001<$ p-value $<0.05$.
Datasets: Conversation Chronicles (CC)
Dialogue Session: Session 3 Session 4 Session 5 Avg.
[Uncaptioned image]Theanine Theanine’s Winning Percentage
          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
All Dialogue History 69% 85% 83% 67% 82% 79% 71% 82% 82% 69% 83% 81%
All Memories and $\mathcal{D}$ 66% 62% 68% 66% 69% 70% 69% 71% 73% 67% 68% 70%
     + Memory Update 54%* 59% 62% 55%* 58% 59% 55%* 63% 64% 54% 60% 62%
Memory Retrieval 67% 64% 71% 70% 67% 68% 71% 66% 71% 69% 66% 70%
     + Memory Update 57% 59% 61% 55%* 57% 62% 57% 62% 66% 56% 60% 63%
RSum-LLM 31% 64% 69% 29% 60% 67% 32% 60% 68% 31% 61% 68%
MemoChat 56%* 73% 73% 57% 69% 71% 55%* 70% 70% 56% 70% 71%
COMEDY 59% 44% 73% 57% 46% 71% 55%* 46% 71% 57% 46% 72%
Table 5: Winning percentage of Theanine in head-to-head G-Eval (of all generated responses) with baselines. Tested with Conversation Chronicles (CC). Color gray shows where ours loses more. Underline: p-value $<0.001$. *: $0.001<$ p-value $<0.05$.
Figure 9: Interface for human evaluation regarding memory linking.
Figure 10: Interface for human evaluation regarding timeline refinement.
Figure 11: Interface for human evaluation regarding referencing past conversations in responses.
Figure 12: Examples of Relation-aware Memory Linking - 1.
Figure 13: Examples of Relation-aware Memory Linking - 2.
Figure 14: Examples of Relation-aware Memory Linking - 3.
Figure 15: Examples of Timeline Refinement and Response Generation.
Figure 16: The prompt for the Relation-aware memory linking.
Figure 17: The prompt for the context-aware timeline refinement.
Figure 18: The prompt for the timeline-augmented response generation.
Figure 19: The prompt for the memory updating mechanism in baselines (i.e., + Memory Update).
Figure 20: The prompt for the G-Eval: Specificity.
Figure 21: The prompt for the G-Eval: Memorability.
Figure 22: The prompt for the G-Eval: Consistency.
Figure 23: The prompt for generating counterfactual QA in TeaFarm.
Figure 24: The prompt for generating session 6 in TeaFarm.
Figure 25: The prompt for evaluating model response in TeaFarm.

### v2

@@
-Seo Hyun Kim1     Kai Tzu-iunn Ong1 Taeyoon Kwon1     Namyoung Kim1     Keummin Ka1     SeongHyeon Bae1 Yohan Jo2     Seung-won Hwang2     Dongha Lee1     Jinyoung Yeo1
+Towards Lifelong Dialogue Agents via Relation-aware Memory Construction and Timeline-augmented Response Generation
 Abstract
-Large language models (LLMs) are capable of processing lengthy dialogue histories during prolonged interaction with users without additional memory modules; however, their responses tend to overlook or incorrectly recall information from the past.
-In this paper, we revisit memory-augmented response generation in the era of LLMs.
-While prior work focuses on getting rid of outdated memories, we argue that such memories can provide contextual cues that help dialogue systems understand the development of past events and, therefore, benefit response generation.
-We present [Uncaptioned image]Theanine, a framework that augments LLMs’ response generation with memory timelines – series of memories that demonstrate the development and causality of relevant past events.
-Along with Theanine, we introduce TeaFarm, a counterfactual-driven question-answering pipeline addressing the limitation of G-Eval in long-term conversations.
-Supplementary videos of our methods and the [Uncaptioned image]TeaBag dataset for TeaFarm evaluation are in https://theanine-693b0.web.app/.
-[Uncaptioned image]Theanine: Revisiting Memory Management in Long-term Conversations with Timeline-augmented Response Generation
-Seo Hyun Kim1thanks: These two authors are the co-first authors and both have the right to list their names first on their CVs.
-Kai Tzu-iunn Ong1 Taeyoon Kwon1thanks: T Kwon, N Kim, and K Ka are the co-second authors.
-Namyoung Kim1     Keummin Ka1     SeongHyeon Bae1 Yohan Jo2     Seung-won Hwang2     Dongha Lee1     Jinyoung Yeo1 1Yonsei University, 2Seoul National University
+To achieve lifelong human-agent interaction, dialogue agents need to constantly memorize perceived information and properly retrieve it for response generation (RG).
+While prior work focuses on getting rid of outdated memories to improve retrieval quality, we argue that such memories provide rich, important contextual cues for RG (e.g., changes in user behaviors) in long-term conversations.
+We present [Uncaptioned image]Theanine, a framework for LLM-based lifelong dialogue agents.
+Theanine discards memory removal and manages large-scale memories by linking them based on their temporal and cause-effect relation.
+Enabled by this linking structure, Theanine augments RG with memory timelines - series of memories representing the evolution or causality of relevant past events.
+Along with Theanine, we introduce TeaFarm, a counterfactual-driven evaluation scheme, addressing the limitation of G-Eval and human efforts in measuring memory-augmented dialogue agents.
+A supplementary video for Theanine and data for TeaFarm are at https://huggingface.co/spaces/ResearcherScholar/Theanine.
+Towards Lifelong Dialogue Agents via Relation-aware Memory Construction and Timeline-augmented Response Generation
+Kai Tzu-iunn Ong1thanks: KT Ong and N Kim are the co-first authors.
+Namyoung Kim1    Minju Gwak1    Hyungjoo Chae1 Taeyoon Kwon1    Yohan Jo2    Seung-won Hwang2    Dongha Lee1    Jinyoung Yeo1 1Yonsei University, 2Seoul National University
 1 Introduction
-Recently, the impressive performance of large language models (LLMs) has caused a trend of having longer conversations with them across multiple sessions (Fui-Hoon Nah et al., 2023).
-Such prolonged interactions require the models to memorize previous events or speaker information, tailoring the responses accordingly to keep up user interest (Adiwardana et al., 2020).
-While the large context windows of LLMs [^1] Recent LLMs, such as GPT-4 (Achiam et al., 2023) and LLama3 (MetaAI, 2024), have contexts windows of 128,000 and 1,040,000 tokens, respectively.
-allow them to process the entire dialogue history for response generation (RG), this often leads to biased attention toward the latest user utterance, ignoring relevant contexts from the past (Liu et al., 2024), as presented in Figure 1 (a).
-Figure 1: Empirical examples of failed responses due to (a) bias to the latest input (b) the absence of an important past event (“afraid of cruise ships”) on the timeline.
-(c) shows a response augmented with the memory timeline.
+Autonomous agents based on large language models (LLMs) have made great progress in various domains, including response generation (Zhou et al., 2023; Kwon et al., 2024; Tseng et al., 2024), where agents ought to constantly keep track of both old and newly introduced information shared with users throughout their service lives (Irfan et al., 2024) and converse accordingly.
+To facilitate such lifelong interaction, studies have proposed to improve dialogue agents’ ability to memorize and correctly recall past information (e.g., discussed topics) in long-term, multi-session conversations.
+Figure 1: Empirical examples of failed responses due to (a) absence of an important past event (“afraid of cruise ships”) on the timeline and (b) bias to the latest input.
+(c) is a response augmented with the memory timeline.
 Figure 2: The overview of Refer to captionTheanine.
 Left: Linking new memories to the memory graph after finishing a dialogue session; Right: Memory timeline retrieval, refinement, and response generation in a new dialogue session.
-One straightforward solution is to condense past conversations into summarized memories and retrieve them to augment response generation in later sessions (Xu et al., 2022a, b; Lu et al., 2023), but the growing span of memories can hinder the quality of memory retrieval as dialogue sessions accumulate.
-Although this can be mitigated by updating the older memories (Bae et al., 2022; Zhong et al., 2024; Kim et al., 2024a), we observe severe information loss.
-As illustrated in Figure 1 (b), an earlier memory on the timeline, which contains an important speaker persona (i.e., “afraid of ships”), is removed during memory update (thus can not be retrieved), yielding a suboptimal response.
-Motivated by these, in this paper, we revisit memory-augmented response generation for long-term conversations and address the aforementioned bottlenecks.
-Specifically, we focus on using the timeline built with relevant memories to augment response generation.
-To this end, we present [Uncaptioned image]Theanine, [^2] L-theanine is an amino acid found in green tea that has been linked to memory improvement (Nguyen et al., 2019).
-a framework of timeline-augmented chain-of-thought reasoning for response generation in long-term conversations.
-Inspired by how humans naturally link new memories to existing memories of relevant events and retrieve them as a whole (Bartlett, 1995), Theanine manages memories with a graph structure, i.e., the memory graph, where memories are linked based on how they relate to each other (Phase I).
-Intuitively, this can be done by linking each memory to others that have high text similarity.
-Yet, we find this naive approach not much beneficial to response generation, due to the plain connection between memories (e.g., “this happened $\rightarrow$ that occurred”).
-To resolve this, we leverage an LLM to dynamically link memories based on their temporal and cause-effect commonsense relations (Hwang et al., 2021) via our designed algorithm.
-Next, in memory retrieval for generating a new response (Phase II-1), instead of solely retrieving top-$k$ memories using current conversation as query, we retrieve the whole memory timeline representing the development of relevant events.
-Then, to tackle the discrepancy between off-line memory construction and online deployment (Phase II-2), Theanine refines the retrieved timeline such that it provides tailored information for the current conversation.
-Finally, in response generation (Phase III), we leverage LLMs’ chain-of-thought (CoT) reasoning ability (Wei et al., 2022) to reason over current conversation and refined timelines, conclude useful information, and generate the final response.
-Our contributions are three-fold:
-  • We revisit memory-augmented response generation in long-term conversations with a novel timeline-augmented framework, [Uncaptioned image]Theanine.
-In both LLM-based (i.e., G-Eval) and human evaluation, Theanine demonstrates (i) better responses that are detailed/non-generic and (ii) efficiency in properly referencing past conversations.
-  • The lack of ground-truth mapping between ongoing conversations (queries) and memories for retrieval poses a challenge in evaluating memory-augmented dialogue systems.
-We present TeaFarm, a counterfactual-driven question-answering pipeline, that assesses systems in referencing past conversations, which does not require human intervention.
-  • We further curate [Uncaptioned image]TeaBag, the counterfactural question-answer data for TeaFarm evaluation on two widely used benchmark datasets of long-term conversations: Multi-Session Chat (MSC) (Xu et al., 2022a) and Conversation Chronicles (CC) (Jang et al., 2023).
+A representative method is to compress past conversations into summarized memories and retrieve them to augment response generation (RG) in later encounters (Xu et al., 2022a; Lu et al., 2023).
+However, the growing span of memories can hinder retrieval quality as conversations accumulate.
+Although it, to some extent, can be solved by updating old memories (Bae et al., 2022; Zhong et al., 2024), this common practice may cause severe information loss.
+As shown in Figure 1 (a), an earlier memory on the timeline, an important persona (“afraid of ships”), is removed during memory update and causes improper RG.
+While using the large context windows of recent LLMs to process all dialogue history/memories is an option to prevent such information loss, [^1] For instance, GPT-4o and Llama 3.1 have context windows of 128K tokens (OpenAI, 2024a; MetaAI, 2024).
+this often leads to biased attention toward the latest user input (Figure 1 (b)), ignoring relevant contexts from the past (Liu et al., 2024).
+These findings highlight two main challenges towards lifelong dialogue agents - (i) Memory construction: how to effectively store large-scale past interactions without removing old memories?
+(ii) Response generation: within the growing memory span, how to identify relevant contextual cues for generating proper responses?
+Motivated by these, we propose to address the above two challenges separately yet complementarily, by (i) discarding memory update to avoid information loss, and preserving relevant memories on the timeline in a linked structure; and (ii) retrieving the timeline as a whole to better catch relevant memories within the growing search span.
+We present [Uncaptioned image]Theanine, [^2] L-theanine is an amino acid found in green tea that has been linked to memory improvement (Nguyen et al., 2019).
+a framework for facilitating lifelong dialogue agents.
+Starting from memory construction (Phase I), instead of stacking raw memory sentences as-is (Xu et al., 2022a), which may affect memory retrieval and also response quality due to the unstructured format of information (Mousavi et al., 2023; Chen et al., 2023), Theanine stores memories in a directed graph.
+In this graph, inspired by how humans naturally link new memories to existing ones of relevant events based on their relation (Bartlett, 1995), memories are linked using their temporal and cause-effect commonsense relations (Hwang et al., 2021).
+Supported by such linking structure, in memory retrieval for RG (Phase II-1), we go beyond conventional top-$k$ retrieval and further obtain the complete timelines to avoid missing out on important memories that have low textual overlap with current conversation (Tao et al., 2023).
+Lastly, to tackle the discrepancy between off-line memory construction and online deployment, Theanine uses an LLM to refine retrieved timelines (Phase II-2) based on current conversation (Madaan et al., 2024), such that they provide tailored information (Chae et al., 2023) for generating responses (Phase III).
+Our contributions are two-fold:
+  • To achieve lifelong dialogue agents, we present [Uncaptioned image]Theanine, an LLM-based framework with a relation-aware memory graph and timeline-augmented response generation for long-term conversations.
+Theanine outperforms baselines across automatic, LLM-based, and human evaluations of RG.
+Also, we confirm that Theanine leads to better retrieval quality, and its intermediate procedures align with human preference.
+To our knowledge, we are the first to model the concept of timelines in memory management and RG.
+  • The lack of golden mapping between conversations and reference memories poses a challenge in assessing memory-augmented agents.
+We present TeaFarm, a counterfactual-driven pipeline evaluating agent performance in referencing the past without human intervention.
 2 Methodologies
-We present [Uncaptioned image]Theanine (Figure 2), a framework that resembles the mechanism of how humans store and retrieve memories for conversations:
+We present [Uncaptioned image]Theanine, a framework for lifelong dialogue agents, inspired by how humans store and retrieve memories for conversations (Figure 2):
 2.1 Memory Graph Construction (Phase I)
-To replicate how humans construct memories, we approach memory management based on a graph structure, i.e., the memory graph $G$:
+To manage large-scale memories and facilitate structured information for RG (Mousavi et al., 2023; Chen et al., 2023), we approach memory management using a memory graph $G$:
 $G=(V,E)$ (1)
@@
 Humans, on the other hand, interpret events by considering the relation between them, such as “how does an event affect the other?” or “why did this person make that change?”.
-Therefore, we adopt a relation-aware memory linking, where each edge between two memories is encoded with their cause-effect commonsense relation $r\in R$, along with their temporal order.
-These relations are adopted from Hwang et al.
-(2021) and modified by us, including HinderedBy, Cause, Want, and 4 more (Appendix A.1).
+Therefore, we adopt a relation-aware memory linking, where an edge between two memories is encoded with their cause-effect commonsense relation $r\in R$, along w/ the temporal order.
+In practice, we adopt the commonly used relations defined by Hwang et al.
+(2021), including HinderedBy, Cause, Want, and 4 more (Appendix B.1).
 We start by determining the relation between $m_{new}$ and each associative memory.
@@
 $M_{a}^{*}=\{m_{i}\in M_{a}\mid\Upsilon(m_{i},m_{new})\in R\}$ (6)
-where $\Upsilon(\cdot,m_{new})\in R$ indicates that the given memory is assigned with an $r\in R$ with $m_{new}$, [^4] Limited by the performance of retrievers, it is possible that an $m\in M_{a}$ does not have a relation, other than just textual overlap, with $m_{new}$.
-We address this by allowing the LLM to output “None”.
+where $\Upsilon(\cdot,m_{new})\in R$ indicates that the given memory is assigned with an $r\in R$ with $m_{new}$, [^4] Limited by retrievers, an $m\in M_{a}$ may not have a relation with $m_{new}$.
+We thus allow the LLM to output “None”.
 and such assigned memories are defined as $M_{a}^{*}$.
-Figure 3: Locating memories to be linked to $m_{new}$.
-We then proceed to link $m_{new}$ to the graph, a straightforward approach is linking $m_{new}$ to all $m\in M_{a}^{*}$.
-Yet, we find it algorithmically inefficient (Figure 7).
-Thus, we selectively link $m_{new}$ to only those $m\in M_{a}^{*}$ that are recent.
-To identify those recent memories, we first locate every connected component $C_{i}\subset G^{t}$ that contains at least one $m\in M_{a}^{*}$, as shown in Figure 3 (a) and (b):
+We then proceed to link $m_{new}$ to the graph. We first locate every connected component $C_{i}\subset G^{t}$ that contains at least one $m\in M_{a}^{*}$, as shown in Figure 3 (a) and (b):
 =$\mathbb{C}=\{C_{i}\subset G^{t}\mid\mathtt{V}(C_{i})\cap M_{a}^{*}\neq\emptyset\ \}$ (7)
 where $\mathbb{C}$ is the collection of those $C$ and $\mathtt{V}(\cdot)$ represents “vertices in”.
-Now, we link $m_{new}$ to the most recent $m\in M_{a}^{*}$ in each $C_{i}\subset\mathbb{C}$ (Figure 3 (c)).
-Therefore, the memories $M_{linked}$ that are linked to $m_{new}$ is defined as follows:
+Then, we link $m_{new}$ to the most recent [^5] Simply linking $m_{new}$ to all $m\in M_{a}^{*}$ costing 25% more API cost for linking without leading to better response.
+$m\in M_{a}^{*}$ in each $C_{i}\subset\mathbb{C}$ (Figure 3 (c)).
+Memories $M_{linked}$ that are linked to $m_{new}$ is defined as follows:
 $M_{linked}=\{\Omega(\mathtt{V}(C_{i})\cap M_{a}^{*})\mid C_{i}\subset\mathbb{C}\}$ (8)
 where $\Omega(\cdot)$ indicates “the most recent memory in”.
-Linking all memories from session $t$ to $G^{t}$, we then obtain a new memory graph $G^{t+1}$.
-The pseudo algorithm for Phase I is in Algorithm 1.
+Figure 3: Locating memories to be linked to $m_{new}$.
+Linking all memories from session $t$ to $G^{t}$, we then obtain a new memory graph $G^{t+1}$. The pseudo algorithm for Phase I is in Algorithm 1.
 2.2 Timeline Retrieval and Timeline Refinement (Phase II)
-Drawing motivation from Figure 1, [Uncaptioned image]Theanine augments response generation (RG) with timelines of relevant events to address the information loss caused by conventional memory management (Xu et al., 2022a; Bae et al., 2022).
-With $G^{t+1}$, Theanine performs these steps for RG in session $t+1$:
+Thanks to the constructed memory graph, Theanine can proceed to augment RG with timelines of relevant events, addressing the information loss in conventional memory management (Figure 1).
+With $G^{t+1}$, Theanine performs the following steps for RG in session $t+1$:
 Preparation: Top-k memory retrieval.
-During the conversation, using the current dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ of $n$ utterances $u$ as query, we first retrieve top-$k$ memories $M_{re}=\{m_{re1},...,m_{rek}\}$.
-Phase II-1: Retrieving raw memory timelines.
-After that, we wish to also access memories that are centered around relevant events.
-Formally, given an $m_{re}\in M_{re}$, we further retrieve the connected component $C_{re}\subset G^{t+1}$ that contains $m_{re}$.
+During the conversation, using the current dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ of $n$ utterances $u$ as query, we retrieve top-$k$ memories $M_{re}=\{m_{re1},...,m_{rek}\}$.
+Phase II-1: Retrieving and untangling raw memory timelines.
+We wish to also access memories centered around $M_{re}$.
+Formally, given $m_{re}\in M_{re}$, we further collect the connected component $C_{re}\subset G^{t+1}$ that contains $m_{re}$ via the linked structure.
 Since this collection of memories (i.e., $C_{re}$) can be “tangled up” together (i.e., connected in a complex manner) due to the graph structure, we proceed to untangle it into several memory timelines, each representing a series of events about $m_{re}$ that starts out similarly yet branches into slightly different development.
@@
 $m_{start}=\Theta(\mathtt{V}(C_{re}))$ (9)
-where $\Theta$ indicates “the oldest memory in”
+where $\Theta$ indicates “the oldest memory in”.
 Figure 4: Extracting raw memory timelines $\tau$ from the retrieved connected component $C_{re}$.
-Next, starting from $m_{start}$, we untangle the memories by tracing through the future direction and extract every possible linear graph containing $m_{re}$ (two in Figure 4) from $C_{re}$, until reaching an endpoint $\tau[-1]$ with an out-degree of 0 (i.e., $deg^{+}(\tau[-1])=0$, which means no directed edge goes out from it).
+Next, starting from $m_{start}$, we untangle memories by tracing through future direction and extract every possible linear graph containing $m_{re}$ (two in Figure 4) from $C_{re}$, until reaching an endpoint $\tau[-1]$ with an out-degree of 0 (i.e., $deg^{+}(\tau[-1])=0$), which means no directed edge goes out from it).
 Each of them is considered a raw memory timeline $\tau$, demonstrating a version of the evolution of $m_{re}$ and its relevant events:
@@
 We then sample $n$ raw timelines $\tau$ from $\mathcal{T}$ [^6] Based on our experiments, we empirically set $n$ to 1.
-Repeating Phase II-1 for all retrieved top-$k$ memories, we collect a set of retrieved raw memory timelines $\mathbb{T}=\cup\,\mathcal{T}$, where $|\mathbb{T}|=k^{*}n$.
+Repeating777“Repeating” is used to explain the algorithm from the perspective of one $m_{re}$.
+In practice, $M_{re}$ are processed together, although processing them 1-by-1 yields the same result.
+Phase II-1 for all retrieved top-$k$ memories, we collect a set of retrieved raw memory timelines $\mathbb{T}=\cup\,\mathcal{T}$, where $|\mathbb{T}|=k^{*}n$.
 Phase II-2: Context-aware timeline refinement.
-Although we have constructed the memory graph using temporal and commonsense relations to improve its informativeness, directly applying the retrieved timelines for response generation can be suboptimal (RQ3, Section 4), because graph construction does not take the current conversation into consideration, i.e., they are constructed off-line.
-In this phase, [Uncaptioned image]Theanine tackles such a discrepancy between off-line memory construction and online deployment (i.e., an ongoing current conversation) via a context-aware timeline refinement.
-Motivated how LLMs can self-refine their previous generation (Madaan et al., 2024).
+Although we have constructed the memory graph using temporal and commonsense relations to improve its informativeness, directly applying retrieved timelines for RG can be suboptimal (RQ3, Section 4), because graph construction does not take current conversation into consideration, i.e., they are constructed off-line.
+In this phase, Theanine tackles such a discrepancy between off-line memory construction and online deployment (i.e., ongoing conversation) via a context-aware timeline refinement.
+Motivated by how LLMs can refine their previous generation (Madaan et al., 2024).
 We leverage LLMs to refine raw timelines into a rich resource of information crafted for the current conversation, by removing redundant information or highlighting information that can come in handy.
-Formally, given the current dialogue $\mathcal{D}$ and the retrieved raw memory timelines $\mathbb{T}$, the LLM tailors all $\tau\in\mathbb{T}$ into refined timelines $\mathbb{T}_{\Phi}$:
+Formally, given the current dialogue $\mathcal{D}$ and retrieved raw timelines $\mathbb{T}$, an LLM tailors $\tau\in\mathbb{T}$ into refined timelines $\mathbb{T}_{\Phi}$:
 $\mathbb{T}_{\Phi}=\{\underset{\tau_{\Phi}}{\text{argmax}}\,P_{\text{LLM}}(\tau_{\Phi}|\mathcal{D},\tau)\mid\tau\in\mathbb{T}\}$ (11)
@@
 2.3 Timeline-augmented Response Generation (Phase III)
-Finally, we move on to response generation.
-Here, [Uncaptioned image]Theanine leverages LLMs’ chain-of-thought (CoT) reasoning to discover user intention from the current conversation (Chae et al., 2023) and selectively extract useful information provided in the refined timelines.
-Formally, given the current dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ and refined timelines $\mathbb{T}_{\Phi}$, the LLM sequentially generates a CoT rationale $\mathcal{R}$ demonstrating how it comprehends $\mathcal{D}$ and $\mathbb{T}_{\Phi}$, and the next response $\bar{u}_{t+1}$:
-¯$\bar{\mathcal{R}}=\underset{\mathcal{R}}{\text{argmax}}\,P_{\text{LLM}}(\mathcal{R}|\mathcal{D},\mathbb{T}_{\Phi})$ (12)
-$\Rightarrow\bar{u}_{n+1}=\underset{u_{n+1}}{\text{argmax}}\,P_{\text{LLM}}({u}_{n+1}|\mathcal{D},\mathbb{T}_{\Phi},\bar{\mathcal{R}})$ (13)
-where $\Rightarrow$ denotes sequential text generation.
+Now, Theanine utilizes the refined timelines for RG. Formally, given $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ and $\mathbb{T}_{\Phi}$, an LLM generates a next response $\bar{u}_{t+1}$:
+$\bar{u}_{n+1}=\underset{u_{n+1}}{\text{argmax}}\,P_{\text{LLM}}({u}_{n+1}|\mathcal{D},\mathbb{T}_{\Phi})$ (12)
 3 Experimental Setups
+Datasets: Multi-session Chat (MSC) Conversation Chronicles (CC)
+Methods / Metrics Bleu-4 Rouge-L Mauve BertScore Bleu-4 Rouge-L Mauve BertScore
+All Dialogue History 1.65 14.89 9.06 86.28 4.90 21.56 26.47 88.13
+All Memories & Current Context $\mathcal{D}$ 1.56 14.89 10.62 86.23 4.41 20.06 38.16 88.02
++ Memory Update (Bae et al., 2022) 1.55 14.77 9.28 86.20 4.34 20.34 34.84 88.03
+Memory Retrieval (Xu et al., 2022a) 1.92 15.49 11.16 86.47 4.93 20.63 33.06 88.07
++ Memory Update (Bae et al., 2022) 1.67 15.30 13.71 86.39 4.46 20.19 34.28 88.02
+Rsum-LLM (Wang et al., 2023) 0.75 11.53 2.45 84.91 0.98 11.42 2.28 85.59
+MemoChat (Lu et al., 2023) 1.42 13.51 7.72 85.96 2.31 15.87 15.12 87.08
+COMEDY (Chen et al., 2024b) 1.06 12.79 7.27 85.29 1.70 13.57 1.95 85.90
+[Uncaptioned image]Theanine (Ours) 1.80 15.37 18.62 86.70 6.85 22.68 64.41 88.58
+Table 1: Automatic evaluation of response quality (average of sessions). Session-specific results are in Appendix F.
 3.1 Datasets of Long-term Conversations
-There are few datasets of long-term conversations.
-First, Multi-Session Chat (MSC) (Xu et al., 2022a), which is built upon the conversations from Persona-Chat (Zhang et al., 2018) by extending their follow-up conversations throughout multiple sessions.
-Soon after MSC, DuLeMon (Xu et al., 2022b) and CareCall (Bae et al., 2022) are proposed for long-term conversations in Mandarin and Korean, respectively.
+There are few datasets of long-term, multi-session conversations.
+First, Multi-Session Chat (MSC) (Xu et al., 2022a), which is built upon Persona-Chat (Zhang et al., 2018) by extending its conversations to multiple (five) sessions.
+Soon after MSC, DuLeMon (Xu et al., 2022b) and CareCall (Bae et al., 2022) are proposed for long-term conversations in Mandarin and Korean.
 Recently, Jang et al.
-(2023) release a new multi-session datasets, Conversation Chronicles (CC).
-Different from MSC, speakers in CC are augmented with diverse relationships, such as “employee and boss”.
-Apart from the above open-domain datasets, there is also Psychological QA, [^6] https://www.xinli001.com/ which addresses long-term conversation in Mandarin regarding clinical scenarios.
-In this work, we adopt MSC and CC for evaluation.
-Since we focus on English conversations, we do not include DuleMon, CareCall, and Psychological QA.
-We leave such multilingual settings and clinical scenarios to future work.
-Datasets: Multi-session Chat (MSC) & Conversation Chronicles (CC)
-Dialogue Session: Session 3 Session 4 Session 5 Avg.
-[Uncaptioned image]Theanine Theanine’s Winning Percentage
-          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
-All Dialogue History 71% 82% 77% 73% 81% 77% 72% 79% 77% 72% 81% 77%
-All Memories and $\mathcal{D}$ 69% 64% 66% 69% 69% 68% 70% 68% 68% 69% 67% 67%
-     + Memory Update 56% 57% 60% 56% 56% 58% 56% 58% 60% 56% 57% 60%
-Memory Retrieval 67% 62% 68% 72% 67% 68% 72% 65% 68% 70% 65% 68%
-     + Memory Update 57% 56% 59% 58% 57% 61% 55% 58% 59% 57% 57% 60%
-RSum-LLM 38% 61% 69% 35% 57% 67% 37% 57% 65% 37% 58% 67%
-MemoChat 60% 74% 71% 60% 69% 72% 60% 72% 71% 60% 72% 71%
-COMEDY 66% 54% 78% 64% 54% 77% 62% 53% 76% 64% 53% 77%
-Golden Responses 63% 76% 61% 63% 73% 61% 62% 72% 61% 63% 73% 61%
-Table 1: Win rates of Theanine in head-to-head G-Eval (of all generated responses) with baselines.
-Spe = specificity; Mem = memorability; Cst = consistency.
-Color green denotes metrics regarding the reference of memories.
-Color gray shows where ours loses more.
-Underline: p-value $<0.001$.
-*: $0.001<$ p-value $<0.05$.
+(2023) release a new dataset, Conversation Chronicles (CC).
+Different from MSC, speakers in CC are augmented with relationships, such as “employee and boss”.
+Apart from the above open-domain datasets, there is also Psychological QA, [^8] https://www.xinli001.com/ which addresses long-term conversations in Mandarin regarding clinical scenarios.
+We use MSC and CC for evaluation. We focus on English conversations, so we do not adopt DuleMon, CareCall, and Psychological QA and leave multilingual and clinical settings to future work.
 3.2 Baselines
-To evaluate Theanine in long-term conversation, besides naive settings using all past dialogues or memories, we include the following baselines.
-Memory retrieval. Following Xu et al. (2022a) and Xu et al. (2022b), this setting uses a retriever to retrieve memories with the current dialogue context to augment response generation.
-+ Memory update. We use LLMs to perform the memory updating mechanism developed by Bae et al. (2022) to update the memory pool at the end of each session to enhance the above baselines.
-RSum-LLM.
-Proposed by Wang et al.
-(2023), it uses an LLM to (i) recursively summarize and update the memory pool at the end of each session, (ii) select top memories based on the current dialogue, and (iii) generate responses.
+To evaluate Theanine, besides naive baselines using all past dialogues or memories, we include:
+Memory Retrieval. Following Xu et al. (2022a), we use a retriever to retrieve memories relevant to the current dialogue context to augment RG.
+Memory Update.
+We use LLMs to run a widely used updating algorithm proposed by Bae et al.
+(2022) at the end of each session.
+The algorithm includes functionalities such as Change, Replace, Delete, Append, and more (see Appendix H).
+RSum-LLM. An LLM-only generative method, which recursively summarizes and updates the memory pool and accordingly generates responses (w/o a retrieval module) (Wang et al., 2023).
 MemoChat.
@@
 Proposed by Chen et al.
-(2024b), rather than memory retrieval, it first summarizes session-level past summaries and then compresses all of them into short sentences of events, user portraits (behavior patterns, emotion, etc.) and user-bot relation, using them to augment response generation.
+(2024b), it uses LLMs to summarize session-level memories, compresses all of them into short events, user portraits (behavioral patterns, emotion, etc.) and user-bot relation, and selects them to augment RG.
 3.3 Models and Implementation Details
 Large language models.
-In all experiments including baselines, we adopt ChatGPT (gpt-3.5-turbo-0125) from OpenAI (2023) for (i) memory summarization, (ii) memory update, and (iii) response generation.
+In all experiments including baselines, we adopt gpt-3.5-turbo-0125 (OpenAI, 2023) for (i) memory summarization (Table 4), (ii) memory update, and (iii) response generation.
 We set the temperature to $0.75$.
 Retrievers.
-We use text-embedding-3-small (OpenAI, 2024b) from OpenAI to calculate the similarity of texts.
-For the identification of top-$j$ associative memories (Phase I-1) and top-$k$ memory retrieval (Phase II Preparation), we set $j$ and $k$ to 3.
-For the “Memory Retrieval” baseline, we retrieve 6 memories following Xu et al.
+We use text-embedding-3-small (OpenAI, 2024b) to calculate text similarity.
+In the identification of top-$j$ associative memories (Phase I-1) and top-$k$ memory retrieval (Phase II), we set $j$ and $k$ to 3.
+For “Memory Retrieval” baseline, we retrieve 6 memories following Xu et al.
 (2022a).
-Dialogue sessions. We perform evaluation using sessions 3-5 from both datasets, as all settings are almost identical before the end of session 2 (there is no memory to update before then).
-4 Evaluation Scheme 1: Machine and Human Evaluation
-To evaluate [Uncaptioned image]Theanine, we first start with G-Eval (Liu et al., 2023), an LLM-based framework for assessing LMs’ responses, and human evaluation.
-[^7] Details for G-Eval and human evaluation are in Appendix D and F.
-Dataset-specific results are in Appendix E.
-Here, we focus on the following research questions (RQs):
-RQ1. Can Theanine generate context-rich responses, rather than those generic and plain?
-RQ2. Can Theanine reference past conversations better than other methods?
-RQ3. Do phases in Theanine make sense?
-RQ4. Do humans agree with Theanine’s procedures, and how do they think of the final responses?
-[Uncaptioned image]Theanine yields responses that are more detail-rich (RQ1).
-In Table 1, we first assess whether generated responses are detail-rich, rather than generic and plain, i.e., specificity (Spe).
-In both datasets, Theanine yields significantly high win rates in the head-to-head comparison with all baselines except for RSum-LLM.
-We presume it is due to the fact that RSum-LLM tends to generate much longer responses (avg.
-77 words per response) than other settings (ours = 32 words; avg of other baselines = 30 words).
-[Uncaptioned image]Theanine utilizes past memories efficiently and correctly (RQ2).
-In long-term conversations, it is crucial and necessary to further assess (i) whether the dialogue system utilizes information from the past when necessary (i.e., memorability) and make sure (ii) the response does not contradict previous conversations (i.e., consistency).
-We present the results in Table 1.
-Theanine demonstrates high win rates in the comparisons with baselines.
-However, COMEDY performs almost as well as us regarding memorability, which we assume is because it always applies two additional LLMs to extract and compress past conversations into concise memos before each response generation (no retrieval).
-However, as we expected, always referencing concise and recent user information can lead to contradiction to older information, thus clearly beaten by Theanine in terms of consistency (Cst).
-Ablation results: Phases in [Uncaptioned image]Theanine matter (RQ3).
-Table 2 presents the results where key phases in Theanine are removed.
-We observe that the removal of relation-aware memory linking (Phase I-2) and context-aware timeline refinement (Phase II-2) leads to large performance drops regarding the reference of past information (i.e., the original Theanine shows higher win rates), especially when timeline refinement is absent.
-This justifies our methodological designs: (i) In graph construction, linking memories with cause-effect relation benefits RG more than naive connection using only text similarity and chronological order; (ii) By further refining memories built off-line (i.e., without considering current conversations) before RG, we can elicit contextual cues that are tailored for the current conversation.
-Multi-session Chat (Session 3-5)
-[Uncaptioned image]Theanine Ours’ Win Rate
-          vs. Mem Cst Avg.
-w/o Relation-aware Linking 62% 65% 64%
-w/o Timeline Refinement 62% 69% 65%
-Conversation Chronicles (Session 3-5)
-w/o Relation-aware Linking 60% 65% 63%
-w/o Timeline Refinement 64% 68% 66%
-Table 2: Ablation results (p-values $<0.001$).
-Figure 5: Results of human evaluation.
-Humans think [Uncaptioned image]Theanine makes sense and properly references past conversations.
-As reported in Figure 5, human judges largely agree (92%) that Theanine’s relation-aware memory linking is performed appropriately, and they all agree (100%) that timeline refinement can elicit more helpful/tailored information for response generation.
-Most importantly, Theanine produces responses that either entail (i.e., faithfully reflect) past conversations (34%) or, at worse, neutral statements (57%) in 91% of the cases from human standpoints, manifesting its’ efficacy in referencing past information in long-term conversations.
-Examples of [Uncaptioned image]Theanine in terms of each phase are in Appendix G.
+Dialogue sessions. We use sessions 3-5 of MSC and CC for evaluations, as all methods are almost the same in sessions 1-2 (no memory to update).
+4 Evaluation Scheme 1: Automatic and Human Evaluations
+To evaluate [Uncaptioned image]Theanine’s responses in long-term conversations, we follow common practices and conduct 3 types of evaluations: (i) Automatic evaluations; (ii) G-Eval (Liu et al., 2023), an LLM-based framework commonly used to evaluate LMs’ generation; (iii) human evaluation.
+We now present the main findings (details, prompts, and interfaces of evaluations in Scheme 1 are in Appendix E):
+(Finding 1) Theanine outperforms baselines in response generation.
+Table 1 reports agent performance in RG regarding both overlap-based and embedding-based metrics: Bleu-4 (Papineni et al., 2002), Rouge-L (Lin, 2004), Mauve (Pillutla et al., 2021), and BertScore (Zhang et al., 2020).
+Theanine, in both datasets, achieves superior response quality than diverse baselines.
+Despite slightly lower scores compared to Memory Retrieval in overlap-based metrics (i.e., B-4 and R-L) in MSC, Theanine largely outperforms it in terms of embedding-based metrics.
+Interestingly, including ours, methods w/o memory update generally yield higher scores, justifying our proposal towards an update-free memory management for lifelong dialogue agents.
+Settings / Metrics B-4 R-L Mauve Bert
+[Uncaptioned image]Theanine (Ours) 4.32 19.03 41.52 87.64
+w/o Relation-aware Linking 4.07 18.58 39.69 87.57
+w/o Timeline Refinement 4.03 18.82 41.34 87.66
+Broken Down, Shuffled Timeline 4.15 18.70 38.49 87.61
+Memory Retrieval 3.43 18.06 22.11 87.27
+Table 2: Performance of our ablations (avg. of datasets).
+(Finding 2 & 3) All phases contribute to performance; retrieving the timeline as a whole brings large improvement over conventional retrieval.
+To gain deeper insights into our design, we show Theanine’s performance when relation-awareness in linking (Phase I-2) and Timeline Refinement (Phase II-2) are ablated.
+Also, to objectively assess whether Theanine’s retrieval (i.e., retrieving the timeline as a whole) improves retrieval quality, we include a setting where retrieved timelines are broken down into randomly ordered events such that retrieved memories during RG are in the same format as conventional top-$k$ retrieval.
+In Table 2, we observe a ranking regarding contribution to performance: relation-aware linking $>$ retrieving timeline as a whole $>$ timeline refinement.
+We thus confirm the efficacy of constructing a memory graph with causal relations.
+Moreover, utilizing such a structure to collect timelines of relevant events yields better RG than conventional retrieval, despite the smaller $k$ (3 vs.
+6) in initial retrieval.
+Refining timelines shows smaller performance gains, suggesting room for improvement in applying them for RG.
+We leave it to future work.
+(Finding 4) Humans and G-Eval both reveal that Theanine retrieves memories that are more helpful for response generation.
+Beyond agent responses, we further investigate how different memory construction methods affect retrieved memories’ helpfulness in augmenting RG.
+Given the same current dialogues as queries for retrieval, Figure 5 (left; G-Eval; right: Human Eval) shows head-to-head comparisons (ours vs baselines) regarding whose retrieved memories benefit response generation more.
+We observe higher win rates of Theanine in all comparisons, especially in human evaluations, suggesting that ours can facilitate a more helpful memory-augmentation for RG.
+Figure 5: Human- (right) and machine-based (left) head-to-head comparisons between ours and baselines regarding the helpfulness of retrieved memories.
+(Finding 5) Humans confirm that Theanine yields responses better entailing past interactions.
+Now that the helpfulness of Theanine’s retrieved memories is validated, we proceed to investigate whether such helpful memories can help us facilitate reliable lifelong dialogue agents.
+For that, we further ask a group of workers to specifically judge whether agent responses entail, contradict, or are neutral to the past via majority voting.
+In Figure 6, we find Theanine not only leading to a small number of contradictory responses (4%) but also showing the largest percentage (68%; out of 100) of responses that entail past conversations, significantly outperforming baselines.
+We argue that it is because our timeline-based approach elicits memories better at representing past interactions between speakers, thus leading to responses that more directly align with the past, which is important for dialogue agents to maintain long-term intimacy with users (Adiwardana et al., 2020).
+Also, such entail, non-contradictory nature of Theanine’s responses highlights its potential for applications in specialized domains, such as personalized agents for clinical scenarios, where entailment between agent responses and users’ past information (e.g., electrical health records or previous consulting sessions) is crucial for diagnostic decison-making (Tseng et al., 2024).
+Figure 6: Human evaluations regarding to what extent the agent responses entail past conversations.
+As a side note, Memory Update yields fewer contradictory responses (2%), indicating a potential trade-off between (i) removing outdated memories to prevent contradiction and (ii) preserving them to get richer information for RG (Kim et al., 2024a).
+(Finding 6) Humans agree with Theanine’s intermediate procedures.
+As reported in Figure 7, judges largely agree (92%) that Theanine properly assigns cause-effect relations to linked memories, explaining its contribution to performance.
+Also, they agree that timeline refinement successfully elicits more helpful information (100%; 100 samples in total) for RG.
+Examples of Theanine’s phases and RG are in Appendix G.
+Figure 7: Human evaluation of our intermediate phases.
 5 Evaluation Scheme 2: TeaFarm – a Counterfactual-driven Evaluation Pipeline for Long-term Conversations
-Despite the success of G-Eval (Liu et al., 2023) in evaluating responses of LMs, accessing dialogue systems’ ability to utilize past memories in long-term conversation still remains challenging.
-This is due to the unavailability of ground-truth mapping between current conversations and the correct memories for retrieval.
-Although we can resolve this by feeding the evaluator LLM (e.g., GPT-4) the entire memory pool or past history and prompt it to determine whether a response correctly recalls past conversations, the evaluation can be limited by the performance of the evaluator LLM itself.
-To overcome this, along with Theanine, we present TeaFarm, a counterfactual-driven pipeline for evaluating memory-augmented response generation in long-term conversations.
-5.1 Testing Dialogue Systems’ Memory via Counterfactual Questions
-In TeaFarm, we proceed to “trick” the dialogue systems into generating incorrect responses, and the system must correctly reference past conversations to avoid being misled by us.
-Specifically, we talk to the dialogue system while acting as if a non-factual statement is true (thus counterfactual).
+Evaluating memory-augmented agents in long-term conversations is non-trivial due to the unavailability of ground-truth mapping between current conversations and correct memories for retrieval.
+Although we may resort to G-Eval by feeding evaluator LLMs (e.g., GPT-4) the entire past history and prompt it to determine whether a response correctly recalls the past, the evaluation can be largely limited by the performance of the evaluator LLM itself (Kim et al., 2024b).
+To overcome this, along with Theanine, we present TeaFarm, a human-free counterfactual-driven pipeline for evaluating memory-augmented response generation in long-term conversations.
+5.1 Testing Dialogue Agents’ Memory via Counterfactual Questions
+In TeaFarm, we proceed to “trick” dialogue agents into generating incorrect responses, and agents must correctly reference past conversations to avoid being misled by us.
+Specifically, we talk to the dialogue agent while acting as if a non-factual statement is true (thus counterfactual).
 Figure 8 presents some examples of counterfactual questions and the corresponding facts.
 Figure 8: Examples of counterfactual questions.
-In practice, when we want to evaluate a system that has been interacting with the user for multiple sessions, we first (1) collect all past conversations and summarize them session by session.
-Then, we (2) feed the LLM [^8] We apply GPT-4 (gpt-4) with a temperature of $0.75$.
-the collected summaries in chronological order such that it can capture the current stage of each discussed event, e.g., “Speaker B has never been to Japan at this moment”, and (3) generate counterfactual questions from the perspective of both speakers (and the correct answers).
-After that, we (4) kick off a new conversation, chat for a while, then (5) naturally ask the counterfactual question, and (6) assess the correctness of its response.
-Overview illustration and prompts for TeaFarm are in Appendix B and H.
+In practice (Figure 10), when we want to evaluate an agent that has been interacting with the user for sessions, we first (1) collect all past conversations and summarize them session by session.
+Then, we (2) feed a question generator LLM [^9] We apply GPT-4 (gpt-4) with a temperature of $0.75$.
+the collected summaries in chronological order such that it can capture the current stage of each discussed event, e.g., “Speaker B does not own a car”, and (3) generate counterfactual questions from the perspective of both speakers (and the correct answers).
+After that, we (4) kick off (i.e., simulate) a new dialogue session, chat for a while, then (5) naturally ask the counterfactual question, and (6) assess the correctness of its response.
+The overview figure, prompts, and synthesized data for TeaFarm are in Appendix CH, and D, respectively.
 5.2 TeaFarm Results
@@
 COMEDY 0.06 0.18 0.12
-[Uncaptioned image]Theanine 0.17 0.24 0.21
-      w/o Relation-aware Linking 0.06 0.13 0.10
-      w/o Timeline Refinement 0.15 0.16 0.16
-Table 3: TeaFarm results regarding correct/total percentages of model responses. Tested in conversations from MSC and CC with 200 counterfactual questions.
-In Table 3, we present model performance in TeaFarm evaluations.
-Theanine outperforms all baselines, especially in CC.
-Ablations perform worse than original Theanine, again proving the efficacy and necessity of relation-aware linking and timeline refinement.
-Surprisingly, all settings result in low correct rates, qualifying TeaFarm as a proper challenge for stress-testing dialogue systems in long-term conversations.
-Interestingly, baselines that utilize retrievers (just like Theanine) show superior performance than those relying on LLMs for memory selection (i.e., RSum-LLM, MemoChat, and COMEDY).
+[Uncaptioned image]Theanine 0.18 0.24 0.21
+      w/o Relation-aware Linking 0.17 0.20 0.19
+      w/o Timeline Refinement 0.16 0.19 0.18
+Table 3: Success rates (SRs) of correctly recalling the past and not being fooled by the counterfactual questions in TeaFarm (tested with 200 questions).
+In Table 3, Theanine shows higher SR than baselines, especially in CC.
+Ablations perform slightly worse than the original, again proving the efficacy of relation-aware linking and timeline refinement.
+Surprisingly, all settings have low SRs, qualifying TeaFarm as a proper pipeline for stress-testing dialogue agents in long-term conversations.
+Interestingly, baselines using retrievers (same as Theanine) show superior performance than settings only relying on LLMs (i.e., RSum-LLM, MemoChat, and COMEDY).
+This, unexpectedly, supports our efforts in developing a new paradigm of memory management in the era of LLMs.
 [^10] Memory update does not affect Memory Retrieval’s performance.
-We believe it is because the counterfactual questions are made to counter the newest stage of each event.
+We believe it is because counterfactual questions are made to counter the newest stage of each event.
 The removal of older memories thus does not have much impact.
-This pattern, unexpectedly, supports our idea of revisiting memory retrieval in the era of LLMs.
-5.3 Dataset for TeaFarm: [Uncaptioned image]TeaBag
-As a byproduct of TeaFarm, we curate [Uncaptioned image]TeaBag, a dataset for TeaFarm evaluation on MSC and CC.
-TeaBag consists of (1) 100 episodes of original conversations from MSC and CC (session 1-5), (2) follow-up conversations (session 6), where the speaker will naturally converse toward the counterfactual questions, and (3) the corresponding answers.
-TeaBag serve as a useful tool for stress-testing if a system can correctly reference past conversations.
-Dataset details are in Appendix C.
-6 Further Discussions
+6 Further Analyses and Discussions
 Cost efficiency.
-In Phase I-2, we choose to link $m_{new}$ to only the most recent associative memories in each $C_{i}$, rather than all associative memories in $M_{a}$.
-Figure 7 compares Theanine with Theanine-All, where $m_{new}$ is linked to all $m\in M_{a}$.
-We find that although Theanine-All costs 25% more, it is beaten by original Theanine.
-We believe this is because linking $m_{new}$ to all $M_{a}$ leads to more short timelines (e.g., an $m_{new}$ from session 4 is directly linked to an old memory from session 1), which provide less information than timelines built with relevant memories from more sessions.
-Regardless, it still yields lower loss rates than strong baselines.
-An open question here is whether we can reach a balance between cost and performance.
-We presume we can further improve the cost efficiency of Theanine in the long-term by training a memory linker using the collected memory-relation pairs from LLMs via knowledge distillation.
-We leave it to future work.
+A concern of Theanine is the API cost.
+Regardless, we argue that it is competitive when both performance and cost are taken into account.
+Figure 9 plots response quality (Mauve score) against the API cost.
+[^11] Calculated Based on session 5, which involves most memories for management.
+We use Mauve for its stronger correlation with humans over diverse metrics (Pillutla et al., 2021).
+We find Theanine and all ablations not only outperform all baselines but also lie on the Pareto frontier, indicating an efficient cost-performance trade-off.
+This suggests Theanine’s value when performance is prioritized over API costs.
+Figure 9: Cost-performance comparisons.
 Growing span of memories.
-Another inquiry left might be whether the growing span of memory would hinder memory retrieval in Theanine if there ever were hundreds of sessions.
-Although this can be a serious issue for conventional memory retrieval, we argue that it will be much mitigated in Theanine because: (i) We retrieve relevant memories as a whole in the form of timelines.
-This works as a safety net in scenarios where an important memory is missed out in top-$k$ retrieval – it might still be retrieved along with the timeline; (ii) We perform timeline refinements.
-Similarly, this also acts as a second insurance against mis-retrieved memories that are actually not helpful.
-Figure 7: API cost and G-Eval results. Red shows baselines that lose even more than Theanine-All in the head-to-head comparisons with original Refer to captionTheanine.
+Another inquiry left might be whether the growing span of memory will eventually hinder memory retrieval in Theanine if there ever are hundreds of sessions.
+Although this can be a serious issue for conventional memory retrieval, we presume that it will be mitigated to some extent in Theanine because: (i) We retrieve relevant memories as a whole in the form of timelines.
+This serves as a safety net in scenarios where an important memory is missed out in top-$k$ retrieval – it might still be retrieved along with the timeline; (ii) We perform timeline refinements to further adjust retrieved timelines based on current dialogue context such that it provided tailored information for the response generator.
+This acts as a second insurance against sub-optimal retrieval.
 7 Related Work
 Long-term Conversations.
-Since the release of MSC (Xu et al., 2022a), there have been many studies on long-term conversations.
-For instance, Bae et al.
-(2022) train a classifier as the memory updater to handle outdated memories in phone call scenarios.
-Recently, based on LLMs’ outstanding in-context learning ability, Wang et al.
-(2023) leverage LLMs to write, update, and select memories for response generation.
-Apart from LLMs’ power, human behaviors have also fostered methods for long-term conversations.
+Since MSC, there are several studies on long-term conversations: Bae et al.
+(2022) train a classifier to update old memories in phone call scenarios.
+As we enter the era of LLMs, Wang et al.
+(2023) and Li et al.
+(2024) leverage LLMs to write and update memories for RG.
+Apart from LLMs’ power, human behaviors also foster methods in this field.
 For example, Zhong et al.
-(2024) apply humans’ forgetting curve to make memories that have been discussed stay in the memory pool for a longer time.
-Also drawing inspiration from how humans manage memories, our [Uncaptioned image]Theanine, on the other hand, addresses the connection between relevant events in both memory construction and retrieval, improving response quality in long-term conversations, which has never been explored to the best of our knowledge.
-Memory-augmentation for personalized dialogue systems.
-The trend of long-term interaction with dialogue systems promotes the adaptation of them for personalized needs (Chen et al., 2024a, c).
+(2024) apply humans’ forgetting curve to make memories that have been discussed stay in the memory pool longer.
+Recently, Maharana et al.
+(2024) also adopt the concept of timelines.
+However, in their work, a timeline is a fixed, pre-defined series of events (potentially unrelated), which simply serve as a user profile for synthesizing dialogue data.
+By contrast, in our work, a timeline is built with relevant events, which are dynamically linked and retrieved as the conversation goes on, serving as useful cues for response generation.
+Memory-augmentation for personalized dialogue agents.
+The trend of long-term interaction with autonomous agents promotes their adaptation for personalized needs (Chen et al., 2024a, c).
 As a pioneer, Xu et al.
-(2022b) train a persona extractor to create user-based memories in long-term conversations.
-However, training a personalized system for long-term use can be non-trivial due to the lack of data (Tseng et al., 2024).
+(2022b) train a persona extractor to create user-based memories.
+However, training personalized agents for long-term use can be non-trivial due to the lack of data (Tseng et al., 2024).
 As a solution, Kim et al.
-(2024a) apply commonsense models and LLMs to augment existing long-term conversation datasets with high-quality persona sentences; Chen et al.
-(2024b) present a training-free LLM-based framework that extracts user behaviors from past conversations for personalized responses.
-Upon the success of LLMs, [Uncaptioned image]Theanine leverages them to build memory timelines, which represent the development of user personas/behaviors throughout the interaction, providing rich cues for personalized responses.
+(2024a) apply commonsense models and LLMs to augment existing long-term data with high-quality persona sentences; Chen et al.
+(2024b) present a training-free LLM-based framework that extracts user behaviors from conversations for personalized RG.
+Upon the success of LLMs, Theanine leverages them to build memory timelines.
+These timelines represent the development of interactions and lead to responses that better entail speaker information, establishing Theanine’s potential for personalized agents.
 8 Conclusions
-This work revisits memory management in multi-session/long-term conversations in the era of LLMs.
-Following how humans store and retrieve relevant memories as a whole, we present [Uncaptioned image]Theanine, a framework of timeline-augmented response generation, which elicits responses that are detail-rich and, more importantly, able to correctly reference previous conversations when necessary.
-We also propose a novel evaluation pipeline (TeaFarm) that addresses the limitation of G-Eval in long-term conversations as well as the dataset for it ([Uncaptioned image]TeaBag).
-We expect this paper to serve as a groundwork and inspiration for exploring more paradigms of memory construction/retrieval in dialogue systems.
+This paper presents a first-ever timeline-based framework for memory-augmented dialogue agents in long-term conversations.
+Comprehensive evaluations show that [Uncaptioned image]Theanine facilitates more beneficial augmentation, leading to responses that are closer to ground-truths and more aligned with speakers’ past interactions.
+Its effectiveness is further confirmed in TeaFarm, a QA-based pipeline we propose to address the limitation of G-Eval and human efforts in assessing memory-augmented RG.
+We expect our novel approach to serve as a new foundation towards lifelong dialogue agents.
 9 Limitations
-This work has the following limitations: First, the maximum amount of dialogue sessions in this study is limited to five due to the lack of longer open-domain English datasets of multi-session conversations.
-Hence, Theanine’s performance in longer conversations has not been investigated.
-As we mentioned in Section 6, we presume that its’ effectiveness can still hold true to some degree even in such scenarios.
-Yet, we do acknowledge the need to apply additional modules that directly address the growing span of dialogue history/memories.
-One possible way is to introduce the summarize-then-compress paradigm in COMEDY (Chen et al., 2024b) to Theanine, which further compresses session-level summaries into a combined short user/event description.
-Second, retrieving and refining timelines for response generation can be less computationally efficient than conventional memory retrieval, as there might exist scenarios where timelines are not necessary, e.g., when there is no need to reference past conversations.
-We believe this can be mitigated by incorporating an additional selection logic/module to choose between memory retrieval, timeline retrieval, and no retrieval at all.
-Lastly, although we include many recent frameworks as baselines, we failed to compare Theanine with MemoryBank (Zhong et al., 2024), a framework inspired by Ebbinghaus’s forgetting curve.
+First, the amount of dialogue sessions in this study is limited to five due to the lack of longer open-domain English datasets.
+As we mentioned in Section 6, we presume that Theanine’s effectiveness can still hold true to some degree in longer conversations.
+Yet, we do acknowledge the need to apply additional modules that directly address the growing span of dialogue history/memories, such as introducing the summarize-then-compress paradigm in COMEDY (Chen et al., 2024b) to compress session-level summaries into a combined short user/event description.
+Second, although we include many recent frameworks as baselines, we failed to compare Theanine with MemoryBank (Zhong et al., 2024), a framework inspired by Ebbinghaus’s forgetting curve.
 This is because the time intervals between sessions in MSC and CC are either mostly measured in hours or not clearly specified (e.g., “a few months later”), whereas MemoryBank requires precise time intervals in days to apply the forgetting curve.
@@
 However, we remain positive about applying such a mechanism to improve Theanine in our ongoing research.
+Lastly, API-based LLMs may introduce risks such as privacy issues.
+A possible solution is to apply Theanine to small open-source LMs for secure, local usage.
+While there exist challenges in data collection, one may achieve this by (i) collecting synthesized conversations with GPT-generated user profiles, (ii) running Theanine on these data, and (iii) using the outputs of each phase to train student LMs (i.e., distillation from teacher LLMs).
 10 Ethical Statements
@@
 Jinyoung Yeo is the corresponding author (jinyeo@yonsei.ac.kr).
-Appendix A Further Implementation Details
-A.1 Cause-effect Commonsense Relations
+Appendix A Appendix Contents
+  • Appendix B.1: Cause-effect Commonsense Relations.
+  • Appendix B.2: Algorithms for [Uncaptioned image]Theanine.
+  • Appendix B.3: Implementation Details on Computational Experiments.
+  • Appendix C: TeaFarm Evaluation.
+  • Appendix D: The [Uncaptioned image]TeaBag Dataset.
+  • Appendix E Details on Evaluation Scheme 1 (G-Eval and Human Evaluations).
+  • Appendix F: Session-specific Evaluation Results.
+  • Appendix G: Empirical Examples.
+  • Appendix H: Prompts Used in This Work.
+  • Appendix I: Further Analyses and Discussions.
+  • Appendix J: Terms for Use of Artifacts.
+Appendix B Further Implementation Details
+B.1 Cause-effect Commonsense Relations
 We adopt and modify commonsense relations from Hwang et al. (2021) for our relation-aware memory linking. Below is the list of our commonsense relations $R$:
@@
 We address this by allowing the LLM to output None.
-A.2 Algorithms for [Uncaptioned image]Theanine
+B.2 Algorithms for [Uncaptioned image]Theanine
 The pseudo algorithms for Phase I and II are provided in Algorithm 1 and 2.
-A.3 Computational Experiments
+B.3 Implementation Details on Computational Experiments
 All computational experiments in this work are based on OpenAI API (OpenAI, 2024a). Thus, no computing infrastructure is required in this work.
-Appendix B TeaFarm Evaluation
+Appendix C TeaFarm Evaluation
 The overview of TeaFarm is illustrated in Figure 10.
-Appendix C The [Uncaptioned image]TeaBag Dataset
-We curate [Uncaptioned image]TeaBag, a dataset for TeaFarm evaluation on MSC and CC. TeaBag consists of:
+Appendix D The [Uncaptioned image]TeaBag Dataset
+As a byproduct of TeaFarm, we curate [Uncaptioned image]TeaBag, a dataset for TeaFarm evaluation on MSC and CC. TeaBag consists of:
   • 100 episodes of original conversations from Multi-Session Chat and Conversational Chronicles (session 1-5; 50 episodes from each dataset)
   • Two pairs of counterfactual QAs for each episode (200 pairs in total).
-  • Two synthesized session 6 for each episode (thus 200 in total), each of which naturally leads to one of the counterfactual questions.
+  • Two synthesized follow-up conversations (i.e., session 6) for each episode (thus 200 in total), each of which naturally guides the conversation from session 5 towards one of the counterfactual questions.
 This dataset is made with GPT-4. The prompt for generation is in Appendix H. We expect future work to apply TeaBag to stress-test if their dialogue system can correctly reference past conversations.
@@
 Also, we have tried our best to confirm that this dataset does not contain any offensive content.
-Appendix D G-Eval
-G-Eval (Liu et al., 2023) is a framework using LLMs with chain-of-thoughts (CoT) and a form-filling paradigm, to assess the quality of models’ text generation.
-G-Eval with GPT-4 has been shown to generate evaluation results that highly align with human judgement (Liu et al., 2023; Kim et al., 2024b).
-Due to the high API cost for it, we conduct head-to-head comparisons between Theanine and 8 baselines using 100 episodes from MSC and CC (around 3,600 Theanine’s responses being compared in each comparison with one baseline), rather than the whole test sets.
-The prompts for G-Eval are in Figure 2021, and 22.
-We use SciPy to calculate p-values.
-[^10] https://scipy.org/
-Appendix E Dataset-Specific G-Eval Results
-We provide dataset-specific results of G-Eval in Table 4 and 5.
-Appendix F Human Evaluation
+For the overview of data collection, please refer to step 1-4 of TeaFarm (Figure 10).
+Appendix E Details on Evaluation Scheme 1
+We perform evaluations using sessions 3-5 from MSC and CC, as all settings are almost identical before the end of session 2, due to the fact that there is no memory to update before then.
+The test sets of MSC and CC contain over 500 and 20,000 episodes of conversations, where each episode has 5 dialogue sessions, yielding 1.2M turns of responses in total.
+Due to the limited budget for generation (both baselines and ours), when not specified, we sample 50 episodes from each dataset for experiments in this paper (around 3.6K conversational turns in total).
+E.1 G-Eval
+G-Eval (Liu et al., 2023) is a framework using LLMs with chain-of-thoughts (CoT) and a form-filling paradigm to assess the quality of models’ text generation.
+G-Eval with GPT-4 has been shown to generate evaluation results that highly align with human judgement (Liu et al., 2023; Kim et al., 2024b) and thus has been widely applied in many LM-based projects.
+We conduct G-Eval on 5 episodes.
+The prompt for evaluating the helpfulness of retrieved memories is in Figure 23. We use SciPy to calculate p-values. [^12] https://scipy.org/
+E.2 Human Evaluation
 We conduct human evaluation, with workers from Amazon Mechanical Turk (AMT). We construct the following three evaluations:
   • Appropriateness of relation-aware memory linking: In this evaluation, we ask the workers to judge whether they agree that the relation-aware linking is properly done for two given memories.
+The interface provided to AMT workers, which includes detailed instructions for human evaluation, is shown in Figure 11.
   • Helpfulness of context-aware timeline refinement: This evaluation requires the workers to determine if they agree that our context-aware refinement really tailors a raw timeline into a resource of useful information for generating the next response.
-  • The quality of Theanine’s responses: Here, the workers are asked to judge if the responses of Theanine correctly refer to past conversations.
+The interface provided to AMT workers, which includes detailed instructions for human evaluation, is shown in Figure 12.
+  • The quality of responses: Here, the workers are asked to judge if the responses correctly refer to past conversations.
 After reading our responses and past memories, they choose whether the responses entail, contradict, or are neutral to past memories.
 To improve evaluation quality, we use GPT-4 to select responses for this specific evaluation based on past memories, addressing the fact that not every turn in the conversation requires previous information to generate the next response (In the other two evaluations, the samples are randomly selected).
+The interface provided to AMT workers, which includes detailed instructions for human evaluation is shown in Figure 13.
+  • The helpfulness of retrieved memories: Given the same dialogue context, human workers are asked to select a memory that is more helpful for generating a next response from ours’ and a baseline’s retrieval.
+The interface provided to AMT workers, which includes detailed instructions for human evaluation is shown in Figure 14
 Each data sample is judged by 3 different workers, and we report the results based on the majority rule.
 In the third evaluation, when every option (entailment, neutral, contradiction) gets one vote, we consider it neutral (13 samples in total).
-These human evaluations are conducted on 300 samples.
-The interfaces provided to AMT workers, which include detailed instructions for human evaluation, are shown in Figure 910, and 11.
+These human evaluations are conducted on 100 conversational turns.
+Appendix F Session-specific Evaluation Results
+We provide session-specific results for automatic evaluations in Table 5.
 Appendix G Empirical Examples
 We provide several empirical examples of [Uncaptioned image]Theanine.
-Examples of relation-aware memory linking are in Figure 1213, and 14.
+Examples of relation-aware memory linking are in Figure 1516, and 17.
 Examples of utilizing refined timeline for response generation are in Figure 18.
@@
   • Timeline-augmented Response generation (Phase III): Figure 21.
-  • + Memory update (baseline): Figure 19.
+  • Memory Update (baseline): Figure 22.
   • RSum-LLM (baseline): We adopt the original prompt from Wang et al. (2023).
@@
   • COMEDY (baseline): We adopt the original prompt from Chen et al. (2024b).
-  • G-Eval: Figure 20,  21, and  22.
+  • G-Eval: The prompt for evaluating the helpfulness of retrieved memories is in Figure 23.
   • Generating counterfactual QA in TeaFarm: Figure 24.
@@
   • Evaluating model responses in TeaFarm: Figure 26.
-Appendix I Terms for Use of Artifacts
+Appendix I Further Analyses and Discussions
+Memory Summarization.
+At the end of each session, we use ChatGPT (gpt-3.5-turbo-0125) to summarize the conversation into memory sentences.
+We conduct examinations on such conversation-memory summarization using 100 randomly sampled sessions from MSC and CC to make sure the quality of raw memories is acceptable.
+The result is in Table 4.
+Memories that … No Can’t judge Yes
+Contain faulty statements 90% 9% 1%
+Miss important statements 95% 4% 1%
+Table 4: Human evaluation of conversation-to-memory summarization in Theanine.
+Appendix J Terms for Use of Artifacts
 We adopt the MSC and CC datasets from Xu et al.
@@
 Figure 10: The overview of TeaFarm Evaluation.
-Datasets: Multi-session Chat (MSC)
-Dialogue Session: Session 3 Session 4 Session 5 Avg.
-[Uncaptioned image]Theanine Theanine’s Winning Percentage
-          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
-All Dialogue History 74% 79% 72% 78% 80% 74% 74% 76% 71% 75% 78% 72%
-All Memories and $\mathcal{D}$ 71% 66% 64% 72% 69% 67% 70% 65% 63% 72% 66% 65%
-     + Memory Update 59% 54%* 58% 58% 53% 58% 56%* 54%* 57%* 58% 54%* 58%
-Memory Retrieval 68% 61% 65% 73% 67% 67% 72% 64% 65% 71% 64% 66%
-     + Memory Update 56%* 53% 57% 61% 58% 61% 53%* 53% 53% 57% 55% 57%
-RSum-LLM 44% 57% 68% 41% 54%* 67% 43% 53% 63% 42% 55% 66%
-MemoChat 65% 75% 70% 64% 69% 72% 65% 74% 72% 65% 73% 71%
-COMEDY 72% 64% 84% 72% 61% 84% 69% 60% 81% 71% 61% 83%
-Table 4: Winning percentage of Theanine in head-to-head G-Eval (of all generated responses) with baselines.
-Tested with Multi-Session Chat (MSC).
-Color gray shows where ours loses more.
-Underline: p-value $<0.001$.
-*: $0.001<$ p-value $<0.05$.
-Datasets: Conversation Chronicles (CC)
-Dialogue Session: Session 3 Session 4 Session 5 Avg.
-[Uncaptioned image]Theanine Theanine’s Winning Percentage
-          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
-All Dialogue History 69% 85% 83% 67% 82% 79% 71% 82% 82% 69% 83% 81%
-All Memories and $\mathcal{D}$ 66% 62% 68% 66% 69% 70% 69% 71% 73% 67% 68% 70%
-     + Memory Update 54%* 59% 62% 55%* 58% 59% 55%* 63% 64% 54% 60% 62%
-Memory Retrieval 67% 64% 71% 70% 67% 68% 71% 66% 71% 69% 66% 70%
-     + Memory Update 57% 59% 61% 55%* 57% 62% 57% 62% 66% 56% 60% 63%
-RSum-LLM 31% 64% 69% 29% 60% 67% 32% 60% 68% 31% 61% 68%
-MemoChat 56%* 73% 73% 57% 69% 71% 55%* 70% 70% 56% 70% 71%
-COMEDY 59% 44% 73% 57% 46% 71% 55%* 46% 71% 57% 46% 72%
-Table 5: Winning percentage of Theanine in head-to-head G-Eval (of all generated responses) with baselines.
-Tested with Conversation Chronicles (CC).
-Color gray shows where ours loses more.
-Underline: p-value $<0.001$.
-*: $0.001<$ p-value $<0.05$.
+Datasets: Multi-session Chat (MSC) & Conversation Chronicals (CC)
+Session: Session 3 Session 4 Session 5
+Methods / Metrics B-4 R-L Mauve Bert B-4 R-L Mauve Bert B-4 R-L Mauve Bert
+All Dialogue History 3.13 18.04 17.34 87.17 3.17 17.96 18.54 87.12 3.53 18.69 17.42 87.31
+All Memories & Current Context $\mathcal{D}$ 2.69 17.29 28.30 87.10 3.10 17.38 22.52 87.06 3.16 17.75 22.35 87.21
++ Memory Update (Bae et al., 2022) 2.80 17.51 22.92 87.11 2.88 17.24 21.22 86.99 3.16 17.90 22.04 87.24
+Memory Retrieval (Xu et al., 2022a) 3.44 18.33 24.68 87.30 3.38 17.55 21.95 87.17 3.46 18.31 19.70 87.33
++ Memory Update (Bae et al., 2022) 3.10 18.08 25.02 87.24 2.99 17.37 25.97 87.10 3.11 17.78 20.99 87.28
+Rsum-LLM (Wang et al., 2023) 0.83 11.30 2.45 85.25 0.87 11.35 2.32 82.20 0.90 11.78 2.33 85.30
+MemoChat (Lu et al., 2023) 1.88 14.83 14.56 86.57 1.81 14.27 10.57 86.43 1.91 14.96 9.13 86.56
+COMEDY (Chen et al., 2024b) 1.14 12.80 4.74 85.53 1.57 13.18 5.16 85.56 1.42 13.56 3.94 85.69
+[Uncaptioned image]Theanine (Ours) 4.21 19.21 45.53 87.70 4.42 18.63 37.84 87.52 4.34 19.23 41.18 87.70
+Table 5: Session-specific results of agent performance in response generation.
 Figure 11: Interface for human evaluation regarding memory linking.
@@
 Figure 13: Interface for human evaluation regarding referencing past conversations in responses.
+Figure 14: Interface for human evaluation regarding the helpfulness of retrieved memories.
 Figure 15: Examples of Relation-aware Memory Linking - 1.
@@
 Figure 22: The prompt for the memory updating mechanism in baselines (i.e., + Memory Update).
-Figure 20: The prompt for the G-Eval: Specificity.
-Figure 21: The prompt for the G-Eval: Memorability.
-Figure 22: The prompt for the G-Eval: Consistency.
+Figure 23: The prompt for the G-Eval: Helpfulness of Retrieved Memories.
 Figure 24: The prompt for generating counterfactual QA in TeaFarm.

### v3

@@
-Towards Lifelong Dialogue Agents via Relation-aware Memory Construction and Timeline-augmented Response Generation
+Towards Lifelong Dialogue Agents via Timeline-based Memory Management
 Abstract
 To achieve lifelong human-agent interaction, dialogue agents need to constantly memorize perceived information and properly retrieve it for response generation (RG).
-While prior work focuses on getting rid of outdated memories to improve retrieval quality, we argue that such memories provide rich, important contextual cues for RG (e.g., changes in user behaviors) in long-term conversations.
+While prior studies focus on getting rid of outdated memories to improve retrieval quality, we argue that such memories provide rich, important contextual cues for RG (e.g., changes in user behaviors) in long-term conversations.
 We present [Uncaptioned image]Theanine, a framework for LLM-based lifelong dialogue agents.
@@
 Enabled by this linking structure, Theanine augments RG with memory timelines - series of memories representing the evolution or causality of relevant past events.
-Along with Theanine, we introduce TeaFarm, a counterfactual-driven evaluation scheme, addressing the limitation of G-Eval and human efforts in measuring memory-augmented dialogue agents.
+Along with Theanine, we introduce TeaFarm, a counterfactual-driven evaluation scheme, addressing the limitation of G-Eval and human efforts when assessing agent performance in integrating past memories into RG.
 A supplementary video for Theanine and data for TeaFarm are at https://huggingface.co/spaces/ResearcherScholar/Theanine.
-Towards Lifelong Dialogue Agents via Relation-aware Memory Construction and Timeline-augmented Response Generation
+Towards Lifelong Dialogue Agents via Timeline-based
+Memory Management
 Kai Tzu-iunn Ong1thanks: KT Ong and N Kim are the co-first authors.
-Namyoung Kim1    Minju Gwak1    Hyungjoo Chae1 Taeyoon Kwon1    Yohan Jo2    Seung-won Hwang2    Dongha Lee1    Jinyoung Yeo1 1Yonsei University, 2Seoul National University
+Namyoung Kim1    Minju Gwak1    Hyungjoo Chae1 Taeyoon Kwon1    Yohan Jo2    Seung-won Hwang2    Dongha Lee1    Jinyoung Yeo1 1Yonsei University, 2Seoul National University {ktio89, namyoung.kim, jinyeo}@yonsei.ac.kr
 1 Introduction
-Autonomous agents based on large language models (LLMs) have made great progress in various domains, including response generation (Zhou et al., 2023; Kwon et al., 2024; Tseng et al., 2024), where agents ought to constantly keep track of both old and newly introduced information shared with users throughout their service lives (Irfan et al., 2024) and converse accordingly.
-To facilitate such lifelong interaction, studies have proposed to improve dialogue agents’ ability to memorize and correctly recall past information (e.g., discussed topics) in long-term, multi-session conversations.
+Autonomous agents based on large language models (LLMs) have made significant progress in various domains, including response generation (Chae et al., 2024; Kwon et al., 2024; Tseng et al., 2024), where agents ought to constantly keep track of both old and newly introduced information shared with users throughout their service lives (Irfan et al., 2024) and converse accordingly.
+To facilitate such lifelong interaction, studies have proposed enhancing dialogue agents’ ability to memorize and accurately recall past information (e.g., discussed topics) in long-term, multi-session conversations.
 Figure 1: Empirical examples of failed responses due to (a) absence of an important past event (“afraid of cruise ships”) on the timeline and (b) bias to the latest input.
@@
 Left: Linking new memories to the memory graph after finishing a dialogue session; Right: Memory timeline retrieval, refinement, and response generation in a new dialogue session.
-A representative method is to compress past conversations into summarized memories and retrieve them to augment response generation (RG) in later encounters (Xu et al., 2022a; Lu et al., 2023).
+A representative approach is to compress past conversations into summarized memories and retrieve them to augment response generation (RG) in later encounters (Xu et al., 2022a; Lu et al., 2023).
 However, the growing span of memories can hinder retrieval quality as conversations accumulate.
-Although it, to some extent, can be solved by updating old memories (Bae et al., 2022; Zhong et al., 2024), this common practice may cause severe information loss.
-As shown in Figure 1 (a), an earlier memory on the timeline, an important persona (“afraid of ships”), is removed during memory update and causes improper RG.
+Although it, to some extent, can be solved by updating old memories (Bae et al., 2022; Zhong et al., 2024), such common practice may cause severe information loss.
+As shown in Figure 1 (a), an earlier memory on the timeline, an important persona (“afraid of ships”), is removed during memory update, resulting in improper RG.
 While using the large context windows of recent LLMs to process all dialogue history/memories is an option to prevent such information loss, [^1] For instance, GPT-4o and Llama 3.1 have context windows of 128K tokens (OpenAI, 2024a; MetaAI, 2024).
 this often leads to biased attention toward the latest user input (Figure 1 (b)), ignoring relevant contexts from the past (Liu et al., 2024).
-These findings highlight two main challenges towards lifelong dialogue agents - (i) Memory construction: how to effectively store large-scale past interactions without removing old memories?
+These findings highlight two main challenges towards lifelong dialogue agents - (i) Memory construction: how to store large-scale past interactions effectively without removing old memories?
 (ii) Response generation: within the growing memory span, how to identify relevant contextual cues for generating proper responses?
-Motivated by these, we propose to address the above two challenges separately yet complementarily, by (i) discarding memory update to avoid information loss, and preserving relevant memories on the timeline in a linked structure; and (ii) retrieving the timeline as a whole to better catch relevant memories within the growing search span.
+Motivated by these, we propose addressing the above two challenges separately yet complementarily, by (i) discarding memory update to avoid information loss, and preserving relevant memories on the timeline in a linked structure; and (ii) retrieving the timeline as a whole to better catch relevant memories within the growing search span.
 We present [Uncaptioned image]Theanine, [^2] L-theanine is an amino acid found in green tea that has been linked to memory improvement (Nguyen et al., 2019).
@@
 Supported by such linking structure, in memory retrieval for RG (Phase II-1), we go beyond conventional top-$k$ retrieval and further obtain the complete timelines to avoid missing out on important memories that have low textual overlap with current conversation (Tao et al., 2023).
-Lastly, to tackle the discrepancy between off-line memory construction and online deployment, Theanine uses an LLM to refine retrieved timelines (Phase II-2) based on current conversation (Madaan et al., 2024), such that they provide tailored information (Chae et al., 2023) for generating responses (Phase III).
+Lastly, to tackle the discrepancy between off-line memory construction and online deployment, Theanine uses an LLM to refine retrieved timelines (Phase II-2) based on current conversation, such that they provide tailored information (Chae et al., 2023) for RG (Phase III).
 Our contributions are two-fold:
-  • To achieve lifelong dialogue agents, we present [Uncaptioned image]Theanine, an LLM-based framework with a relation-aware memory graph and timeline-augmented response generation for long-term conversations.
-Theanine outperforms baselines across automatic, LLM-based, and human evaluations of RG.
-Also, we confirm that Theanine leads to better retrieval quality, and its intermediate procedures align with human preference.
-To our knowledge, we are the first to model the concept of timelines in memory management and RG.
+  • To achieve lifelong dialogue agents, we present [Uncaptioned image]Theanine, an LLM-based framework with a relation-aware memory graph and timeline augmentation for long-term conversations.
+Theanine outperforms representative baselines across automatic, LLM-based, and human evaluations of RG.
+Also, we confirm that Theanine leads to higher retrieval quality, and its procedures align with human preference.
+To our knowledge, we are the first to model the use of timelines (i.e., linked relevant memories) in memory management and response generation.
   • The lack of golden mapping between conversations and reference memories poses a challenge in assessing memory-augmented agents.
@@
 2 Methodologies
-We present [Uncaptioned image]Theanine, a framework for lifelong dialogue agents, inspired by how humans store and retrieve memories for conversations (Figure 2):
+We present [Uncaptioned image]Theanine, a framework for lifelong dialogue agents inspired by how humans store and retrieve memories for conversations (Figure 2):
 2.1 Memory Graph Construction (Phase I)
@@
 }}}m_{start},m_{re}\in\tau\ \\ &\land deg^{+}(\tau[-1])=0\}\end{split}$ (10)
-We then sample $n$ raw timelines $\tau$ from $\mathcal{T}$ [^6] Based on our experiments, we empirically set $n$ to 1.
+We then sample $n$ raw timelines $\tau$ from $\mathcal{T}$ [^6] We empirically set $n$ to 1, as we observe a high degree of overlap across timelines extracted from the same $C_{re}$, which can lead to redundant information (i.e., input tokens) for RG.
 Repeating777“Repeating” is used to explain the algorithm from the perspective of one $m_{re}$.
@@
 [Uncaptioned image]Theanine (Ours) 1.80 15.37 18.62 86.70 6.85 22.68 64.41 88.58
-Table 1: Automatic evaluation of response quality (average of sessions). Session-specific results are in Appendix F.
+Table 1: Automatic evaluation of response quality (average of sessions).
 3.1 Datasets of Long-term Conversations
-There are few datasets of long-term, multi-session conversations.
-First, Multi-Session Chat (MSC) (Xu et al., 2022a), which is built upon Persona-Chat (Zhang et al., 2018) by extending its conversations to multiple (five) sessions.
+There are limited datasets for long-term, multi-session conversations.
+Firstly, Multi-Session Chat (MSC) (Xu et al., 2022a), is built upon Persona-Chat (Zhang et al., 2018) by extending its conversations to multiple (five) sessions.
 Soon after MSC, DuLeMon (Xu et al., 2022b) and CareCall (Bae et al., 2022) are proposed for long-term conversations in Mandarin and Korean.
@@
 (2023) release a new dataset, Conversation Chronicles (CC).
-Different from MSC, speakers in CC are augmented with relationships, such as “employee and boss”.
-Apart from the above open-domain datasets, there is also Psychological QA, [^8] https://www.xinli001.com/ which addresses long-term conversations in Mandarin regarding clinical scenarios.
-We use MSC and CC for evaluation. We focus on English conversations, so we do not adopt DuleMon, CareCall, and Psychological QA and leave multilingual and clinical settings to future work.
+Unlike MSC, CC augments speakers with defined relationships, such as “employee and boss”.
+Apart from these open-domain datasets, the Psychological QA, [^8] https://www.xinli001.com/ addresses long-term conversations under clinical scenarios in Mandarin.
+We opt for MSC and CC for evaluation to focus on English conversations, leaving multilingual and domain-specific conversations (e.g., DuleMon, CareCall, and Psychological QA) to future work.
 3.2 Baselines
-To evaluate Theanine, besides naive baselines using all past dialogues or memories, we include:
+To evaluate Theanine, in addition to naive baselines that utilize all past dialogues or memories, we incorporate the following settings:
 Memory Retrieval. Following Xu et al. (2022a), we use a retriever to retrieve memories relevant to the current dialogue context to augment RG.
 Memory Update.
-We use LLMs to run a widely used updating algorithm proposed by Bae et al.
-(2022) at the end of each session.
-The algorithm includes functionalities such as Change, Replace, Delete, Append, and more (see Appendix H).
-RSum-LLM. An LLM-only generative method, which recursively summarizes and updates the memory pool and accordingly generates responses (w/o a retrieval module) (Wang et al., 2023).
+We utilize LLMs to implement a widely used updating algorithm proposed by Bae et al.
+(2022) at the end of each dialogue session.
+This algorithm includes functionalities such as Change, Replace, Delete, Append, and more (see Appendix H).
+RSum-LLM. An LLM-only generative method that recursively summarizes and updates the memory pool, generating responses w/o a retrieval module (Wang et al., 2023).
 MemoChat.
 Proposed by Lu et al.
-(2023), it uses LLMs’ CoT reasoning ability to (i) conclude important memories from past conversations in a structured topic-summary-dialogue manner, (ii) select memories, and (ii) generate responses.
+(2023), it leverages LLMs’ CoT reasoning ability to (i) conclude important memories from past conversations in a structured topic-summary-dialogue manner, (ii) select memories, and (ii) generate responses.
 COMEDY.
 Proposed by Chen et al.
-(2024b), it uses LLMs to summarize session-level memories, compresses all of them into short events, user portraits (behavioral patterns, emotion, etc.) and user-bot relation, and selects them to augment RG.
+(2024b), it uses LLMs to summarize session-level memories, compresses all of them into short events, user portraits (behavioral patterns, emotion, etc.) and user-bot relation.
+It then selects compressed memories to augment response generation.
 3.3 Models and Implementation Details
 Large language models.
-In all experiments including baselines, we adopt gpt-3.5-turbo-0125 (OpenAI, 2023) for (i) memory summarization (Table 4), (ii) memory update, and (iii) response generation.
-We set the temperature to $0.75$.
+In all experiments, including baselines, we adopt gpt-3.5-turbo-0125 (OpenAI, 2023) for (i) memory summarization (Table 6), (ii) memory update, and (iii) response generation.
+Temperature is set to $0.75$.
 Retrievers.
-We use text-embedding-3-small (OpenAI, 2024b) to calculate text similarity.
+We use text-embedding-3-small (OpenAI, 2024b) to calculate text similarity for settings involving retrievers.
 In the identification of top-$j$ associative memories (Phase I-1) and top-$k$ memory retrieval (Phase II), we set $j$ and $k$ to 3.
-For “Memory Retrieval” baseline, we retrieve 6 memories following Xu et al.
+For the “Memory Retrieval” baseline, we set $k=6$ following Xu et al.
 (2022a).
-Dialogue sessions. We use sessions 3-5 of MSC and CC for evaluations, as all methods are almost the same in sessions 1-2 (no memory to update).
+Dialogue sessions. We use sessions 3-5 of MSC and CC for evaluations, as all methods are almost identical in session 1 $\sim$ 2 (no memory to update).
 4 Evaluation Scheme 1: Automatic and Human Evaluations
 To evaluate [Uncaptioned image]Theanine’s responses in long-term conversations, we follow common practices and conduct 3 types of evaluations: (i) Automatic evaluations; (ii) G-Eval (Liu et al., 2023), an LLM-based framework commonly used to evaluate LMs’ generation; (iii) human evaluation.
-We now present the main findings (details, prompts, and interfaces of evaluations in Scheme 1 are in Appendix E):
+We now present several key findings (details, prompts, and interfaces of evaluations in Scheme 1 are in Appendix E):
 (Finding 1) Theanine outperforms baselines in response generation.
-Table 1 reports agent performance in RG regarding both overlap-based and embedding-based metrics: Bleu-4 (Papineni et al., 2002), Rouge-L (Lin, 2004), Mauve (Pillutla et al., 2021), and BertScore (Zhang et al., 2020).
-Theanine, in both datasets, achieves superior response quality than diverse baselines.
-Despite slightly lower scores compared to Memory Retrieval in overlap-based metrics (i.e., B-4 and R-L) in MSC, Theanine largely outperforms it in terms of embedding-based metrics.
-Interestingly, including ours, methods w/o memory update generally yield higher scores, justifying our proposal towards an update-free memory management for lifelong dialogue agents.
+Table 1 presents the agent performance in RG regarding both overlap-based and embedding-based metrics: Bleu-4 (Papineni et al., 2002), Rouge-L (Lin, 2004), Mauve (Pillutla et al., 2021), and BertScore (Zhang et al., 2020).
+Across both datasets, Theanine, achieves superior response quality than various baselines.
+Although, compared to Memory Retrieval, Theanine scores slightly lower in overlap-based metrics (i.e., B-4 and R-L) in MSC, it largely outperforms Memory Retrieval in embedding-based metrics.
+Interestingly, including ours, methods without memory update generally yield higher scores, justifying our proposal towards an update-, removal-free memory management for lifelong dialogue agents.
+(Finding 2 & 3) All phases contribute to performance; retrieving the timeline as a whole brings large improvement over conventional retrieval.
+To gain deeper insights into our design, we investigate the impact of removing Theanine’s relation-awareness during memory linking (Phase I-2) and Timeline Refinement (Phase II-2).
+Also, to objectively assess whether Theanine’s retrieval (i.e., retrieving the timeline as a whole) improves retrieval quality, we include a setting where retrieved timelines are broken down into randomly ordered events such that retrieved memories during RG are in the same format as conventional top-$k$ retrieval.
+In Table 2, we observe a ranking in terms of contribution to performance: relation-aware linking $>$ retrieving timeline as a whole $>$ timeline refinement.
+This observation confirms the efficacy of constructing a memory graph with causal relations.
+Moreover, utilizing this graph structure to collect timelines of relevant events yields higher RG quality than conventional retrieval, despite the smaller $k$ (3 vs.
+6) in initial retrieval.
+Refining timelines shows smaller performance gains, suggesting room for improvement in applying them for RG.
+We leave it to future work.
 Settings / Metrics B-4 R-L Mauve Bert
@@
 Table 2: Performance of our ablations (avg. of datasets).
-(Finding 2 & 3) All phases contribute to performance; retrieving the timeline as a whole brings large improvement over conventional retrieval.
-To gain deeper insights into our design, we show Theanine’s performance when relation-awareness in linking (Phase I-2) and Timeline Refinement (Phase II-2) are ablated.
-Also, to objectively assess whether Theanine’s retrieval (i.e., retrieving the timeline as a whole) improves retrieval quality, we include a setting where retrieved timelines are broken down into randomly ordered events such that retrieved memories during RG are in the same format as conventional top-$k$ retrieval.
-In Table 2, we observe a ranking regarding contribution to performance: relation-aware linking $>$ retrieving timeline as a whole $>$ timeline refinement.
-We thus confirm the efficacy of constructing a memory graph with causal relations.
-Moreover, utilizing such a structure to collect timelines of relevant events yields better RG than conventional retrieval, despite the smaller $k$ (3 vs.
-6) in initial retrieval.
-Refining timelines shows smaller performance gains, suggesting room for improvement in applying them for RG.
-We leave it to future work.
-(Finding 4) Humans and G-Eval both reveal that Theanine retrieves memories that are more helpful for response generation.
-Beyond agent responses, we further investigate how different memory construction methods affect retrieved memories’ helpfulness in augmenting RG.
-Given the same current dialogues as queries for retrieval, Figure 5 (left; G-Eval; right: Human Eval) shows head-to-head comparisons (ours vs baselines) regarding whose retrieved memories benefit response generation more.
-We observe higher win rates of Theanine in all comparisons, especially in human evaluations, suggesting that ours can facilitate a more helpful memory-augmentation for RG.
+(Finding 4) Humans and G-Eval reveal that Theanine leads to higher retrieval quality regarding both helpfulness and accuracy.
+Beyond agent responses, we further investigate how different memory construction methods affect the quality of memory retrieval.
+Given the same current dialogues as queries for memory retrieval, Figure 5 shows head-to-head comparisons (ours vs.
+baselines) regarding whose retrieved memories more effectively benefit RG.
+We observe higher win rates for Theanine in all comparisons, especially in human evaluations.
+This suggests that our method can facilitate more helpful memory augmentation for response generation.
+In addition to helpfulness, objectively measuring retrieval accuracy is crucial.
+Since existing datasets of long-term conversations do not provide a golden mapping between dialogue contexts and memories (i.e., golden memories for retrieval), we identify 50 dialogue contexts (i.e., test instances) that require a past memory for RG, and manually measure the retrieval accuracy of different agents.
+The results shown in Table 3 indicate that Theanine and its ablations demonstrate higher retrieval accuracy than baselines, and the ranking here aligns with Table 1 and success rates in Table 4.
 Figure 5: Human- (right) and machine-based (left) head-to-head comparisons between ours and baselines regarding the helpfulness of retrieved memories.
+Methods (Agents) Golden Memory is Retrieved/collected (%)
+Memory Retrieval 68.00
++ Memory Update 64.00
+MemoChat 56.00
+COMEDY 48.00
+[Uncaptioned image]Theanine (Ours) 72.00
+Table 3: Human evaluation of the accuracy of memory retrieval (we examine 50 test instances).
 (Finding 5) Humans confirm that Theanine yields responses better entailing past interactions.
-Now that the helpfulness of Theanine’s retrieved memories is validated, we proceed to investigate whether such helpful memories can help us facilitate reliable lifelong dialogue agents.
+Now that the helpfulness of Theanine’s retrieved memories is validated, we proceed to investigate whether such helpful memories contribute towards reliable lifelong human-agent interaction.
 For that, we further ask a group of workers to specifically judge whether agent responses entail, contradict, or are neutral to the past via majority voting.
-In Figure 6, we find Theanine not only leading to a small number of contradictory responses (4%) but also showing the largest percentage (68%; out of 100) of responses that entail past conversations, significantly outperforming baselines.
-We argue that it is because our timeline-based approach elicits memories better at representing past interactions between speakers, thus leading to responses that more directly align with the past, which is important for dialogue agents to maintain long-term intimacy with users (Adiwardana et al., 2020).
-Also, such entail, non-contradictory nature of Theanine’s responses highlights its potential for applications in specialized domains, such as personalized agents for clinical scenarios, where entailment between agent responses and users’ past information (e.g., electrical health records or previous consulting sessions) is crucial for diagnostic decison-making (Tseng et al., 2024).
+In Figure 6, Theanine not only leads to a small number of contradictory responses (4%) but also demonstrates the largest percentage (68%; out of 100) of responses that entail past conversations, significantly outperforming baselines.
+We argue that it is because our timeline-based approach elicits memories better at representing past interactions between speakers, thus leading to responses more directly aligned with the past.
+This alignment is important for dialogue agents to maintain long-term intimacy with users (Adiwardana et al., 2020).
+Furthermore, such entailing and non-contradictory nature of Theanine’s responses highlights its potential for applications in specialized domains, such as personalized agents for clinical scenarios, where entailment between agent responses and users’ past information (e.g., electrical health records or previous consulting sessions) is crucial for diagnostic decison-making (Tseng et al., 2024).
 Figure 6: Human evaluations regarding to what extent the agent responses entail past conversations.
@@
 (Finding 6) Humans agree with Theanine’s intermediate procedures.
-As reported in Figure 7, judges largely agree (92%) that Theanine properly assigns cause-effect relations to linked memories, explaining its contribution to performance.
+As reported in Figure 7, judges largely agree (92%) that Theanine properly assigns cause-effect relations to linked memories, which explains its contribution to performance.
 Also, they agree that timeline refinement successfully elicits more helpful information (100%; 100 samples in total) for RG.
@@
 The removal of older memories thus does not have much impact.
+To provide insight regarding conversation scenarios that are challenging for dialogue agents, we present case studies of how Theanine fail in TeaFarm in Appendix G.
 6 Further Analyses and Discussions
@@
 [^11] Calculated based on session 5, which involves most memories for management.
-We use Mauve for its stronger correlation with humans over diverse metrics (Pillutla et al., 2021).
+We use Mauve for its stronger correlation with humans (Pillutla et al., 2021).
 We find Theanine and all ablations not only outperform all baselines but also lie on the Pareto frontier, indicating an efficient cost-performance trade-off.
 This suggests Theanine’s value when performance is prioritized over API costs.
+Actual API costs and results based on B-4, R-L, and Bert scores are in Appendix I.
 Figure 9: Cost-performance comparisons.
+Time efficiency.
+Time efficiency can be an important consideration when deploying Theanine to real-world scenarios having richer events.
+Figure 10 shows time-performance comparisons regarding both “memory construction” and “retrieval + RG” also using the Pareto frontier.
+Similarly, Theanine and many of its ablations demonstrate an efficient time-performance trade-off.
+Figure 10: Time-performance comparisons.
+Additional comparison: Memory Retrieval with a dynamically-changing k.
+Due to Theanine’s graph-based procedures, the response generator may access different amounts of memories during RG depending on given contexts (i.e., queries used by the retriever) and when the conversation takes place (i.e., an earlier or a later session), whereas conventional methods (Xu et al., 2022a; Bae et al., 2022) often have a fixed number $k$ of memories retrieved for RG.
+Therefore, to further quantify the effect of our proposed timeline-based management and augmentation, we compare Theanine to Memory Retrieval with a dynamic $k$, where k dynamically changes based on the number of collected memories in Theanine for each specific test data.
+In other words, if Theanine uses timelines to collect $k$ memories during RG for a test instance $\mathcal{D}_{i}$, baselines will also be retrieving $k$ memories for generating a response for $\mathcal{D}_{i}$.
+Methods / Metrics Bleu-4 Rouge-L Mauve Bert
+Memory Retrieval (dynamic k) 3.06 17.97 33.33 87.32
++ Memory Update 2.68 17.19 28.49 87.11
+[Uncaptioned image]Theanine (Ours) 4.22 19.22 45.53 87.70
+Table 5: Additional comparison, where $k$ in Memory Retrieval is dynamically modified for each test instance.
+In Table 5, we can observe that when the number of memories is matched, ours outperforms both baselines despite the same amount of memories being provided.
+We assume this is because: (i) our graph-based retrieval helps us collect more beneficial memories than conventional retrieval; (ii) addressing the relation between events and shaping them based on dialogue contexts can facilitate richer contextual cues for RG.
 Growing span of memories.
-Another inquiry left might be whether the growing span of memory will eventually hinder memory retrieval in Theanine if there ever are hundreds of sessions.
-Although this can be a serious issue for conventional memory retrieval, we presume that it will be mitigated to some extent in Theanine because: (i) We retrieve relevant memories as a whole in the form of timelines.
-This serves as a safety net in scenarios where an important memory is missed out in top-$k$ retrieval – it might still be retrieved along with the timeline; (ii) We perform timeline refinements to further adjust retrieved timelines based on current dialogue context such that it provided tailored information for the response generator.
+Another inquiry is whether the growing span of memory will eventually hinder retrieval in Theanine if there ever are hundreds of sessions.
+Although this may be a serious issue for conventional methods, we presume that it will be partially mitigated in Theanine, as: (i) We retrieve relevant memories as a whole in the form of timelines.
+This serves as a safety net in scenarios where an important memory is missed out in top-$k$ retrieval–it may be collected via the linked structure; (ii) We refine retrieved timelines based on current dialogue such that they provide tailored information for RG.
 This acts as a second insurance against sub-optimal retrieval.
@@
 Long-term conversations.
-Since MSC, there are several studies on long-term conversations: Bae et al.
+Since MSC, there have been several studies on long-term conversations: Bae et al.
 (2022) train a classifier to update old memories in phone call scenarios.
-As we enter the era of LLMs, Wang et al.
-(2023) and Li et al.
-(2024) leverage LLMs to write and update memories for RG.
+As we enter the era of LLMs, Li et al.
+(2024) leverages LLMs to write and update memories for RG.
 Apart from LLMs’ power, human behaviors also foster methods in this field.
 For example, Zhong et al.
-(2024) apply humans’ forgetting curve to make memories that have been discussed stay in the memory pool longer.
-Recently, Maharana et al.
+(2024) apply humans’ forgetting curve to make memories that have been discussed exist longer.
+Recently, Park et al.
+(2023) and Maharana et al.
 (2024) also adopt the concept of timelines.
-However, in their work, a timeline is a fixed, pre-defined series of events (potentially unrelated), which simply serve as a user profile for synthesizing dialogue data.
-By contrast, in our work, a timeline is built with relevant events, which are dynamically linked and retrieved as the conversation goes on, serving as useful cues for response generation.
+However, Park et al.
+(2023) focus on tagging the timestamp (e.g., “22:00”) of events and does not explicitly model the connection between them, and, in Maharana et al.
+(2024), a timeline is a fixed, pre-defined series of events (potentially unrelated) which simply serve as a user profile for synthesizing dialogue data.
+By contrast, in our work, a timeline is built with relevant events, which are dynamically linked based on their causal relations and retrieved as the conversation goes on, benefitting our goal of consistent memory tracking and integration.
 Memory-augmentation for personalized dialogue agents.
@@
 8 Conclusions
-This paper presents a first-ever timeline-based framework for memory-augmented dialogue agents in long-term conversations.
-Comprehensive evaluations show that [Uncaptioned image]Theanine facilitates more beneficial augmentation, leading to responses that are closer to ground-truths and more aligned with speakers’ past interactions.
-Its effectiveness is further confirmed in TeaFarm, a QA-based pipeline we propose to address the limitation of G-Eval and human efforts in assessing memory-augmented RG.
-We expect our novel approach to serve as a new foundation towards lifelong dialogue agents.
-9 Limitations
+This paper presents the first-ever timeline-based memory management and augmentation framework, [Uncaptioned image]Theanine, for autonomous agents in long-term conversations.
+Applying Theanine, we develop a dialogue agent that efficiently addresses the constant, lifelong tracking of memories and their integration for response generation throughout its service life.
+Comprehensive evaluations show that [Uncaptioned image]Theanine can facilitate more beneficial memory augmentation, leading to responses that are closer to ground truths and more aligned with speakers’ past interactions.
+Theanine’s effectiveness is further confirmed in TeaFarm, a counterfactual-driven pipeline we design to address the limitation of G-Eval and human efforts in assessing memory augmentation.
+We expect our novel approaches to serve as a new foundation for future efforts towards lifelong dialogue agents.
+Limitations
 First, the amount of dialogue sessions in this study is limited to five due to the lack of longer open-domain English datasets.
@@
 While there exist challenges in data collection, one may achieve this by (i) collecting synthesized conversations with GPT-generated user profiles, (ii) running Theanine on these data, and (iii) using the outputs of each phase to train student LMs (i.e., distillation from teacher LLMs).
-10 Ethical Statements
-LLMs might generate harmful, biased, offensive, sexual content.
+Ethical Statements
+LLMs might generate harmful, biased, offensive, and sexual content.
 Authors avoid such content from appearing in this paper.
@@
 We ensure an effective pay rate higher than 20\$ per hour based on the estimated time required to complete the tasks.
-11 Acknowledgments
-This work was partly supported by an IITP grant funded by the Korean Government (MSIT) (No.
-RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)).
+Acknowledgments
+This work was mainly supported by STEAM R&D Project, NRF, Korea (RS-2024-00454458) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (No.
+RS-2024-00457882, National AI Research Lab Project), and was partially supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00333484; RS-2024-00414981).
 Jinyoung Yeo is the corresponding author (jinyeo@yonsei.ac.kr).
 Appendix A Appendix Contents
-  • Appendix B.1: Cause-effect Commonsense Relations.
+  • Appendix B.1: Cause-effect Commonsense Relations Adopted.
   • Appendix B.2: Algorithms for [Uncaptioned image]Theanine.
@@
   • Appendix E Details on Evaluation Scheme 1 (G-Eval and Human Evaluations).
-  • Appendix F: Session-specific Evaluation Results.
+  • Appendix F: Session-specific Results of Automatic Evaluation.
   • Appendix G: Empirical Examples.
   • Appendix H: Prompts Used in This Work.
-  • Appendix I: Further Analyses and Discussions.
+  • Appendix I: Further Analyses.
   • Appendix J: Terms for Use of Artifacts.
@@
 Appendix C TeaFarm Evaluation
-The overview of TeaFarm is illustrated in Figure 10.
+The overall pipeline of TeaFarm is illustrated in Figure 11.
 Appendix D The [Uncaptioned image]TeaBag Dataset
@@
 Appendix G Empirical Examples
+Outputs from Theanine.
 We provide several empirical examples of [Uncaptioned image]Theanine.
-Examples of relation-aware memory linking are in Figure 1516, and 17.
+Examples of relation-aware memory linking are in Figure 1617, and 18.
 Examples of utilizing refined timeline for response generation are in Figure 19.
+How Theanine fails in TeaFarm.
+We present failure cases where Theanine fails to pass the TeaFarm test in Figure 20 and Figure 21.
+In Figure 20, although the conversation has shifted to “librarian”, the similarity-based retriever retrieves unhelpful memories due to the huge portion of “kid” in the context.
+While a helpful memory (i.e., “A is a retired libraria”) is eventually caught by our designed timeline structure, the LLM still hallucinate.
+We assume it is due to the noises introduced by those highly-ranked, yet irrelevant memories, and it highlights the need for addressing helpfulness ranking among retrieved memories in lifelong dialogue systems.
+Figure 21 shows a failure case, where Theanine successfully retrieves the correct memories but generates an improper response.
+We hypothesize that this is because relation-aware linking and context-aware timeline refinement may sometimes make the length of input tokens too long such that the agent cannot properly utilize key information provided.
+We believe this can be resolved to an extent via dedicated prompt (i.e., the prompt for RG) engineering.
+We leave this to future work.
 Appendix H Prompts
@@
   • Evaluating model responses in TeaFarm: Figure 29.
-Appendix I Further Analyses and Discussions
+Appendix I Further Analyses
 Memory summarization.
 At the end of each session, we use ChatGPT (gpt-3.5-turbo-0125) to summarize the conversation into memory sentences.
-We conduct examinations on such conversation-memory summarization using 100 randomly sampled sessions from MSC and CC to make sure the quality of raw memories is acceptable.
+We conduct examinations on such summarization using 100 randomly sampled sessions from MSC and CC to make sure the quality of raw memories is acceptable.
 The result is in Table 6.
@@
 Table 6: Human evaluation of conversation-to-memory summarization in Theanine.
+Cost-efficiency trade-off assessed using other metrics.
+In Section 6, we have presented methods having an efficient cost-performance trade-off (i.e., are Pareto-efficient) by plotting the Mauve score against API cost (Figure 9).
+We present methods that are Pareto-efficient when considering the other three metrics used in our study, i.e., B-4, R-L, and Bert Score, in Table 7.
+Agents B-4 R-L Bert Score
+All Dialogue History
+All Memories
++ Update
+Memory Retrieval $\checkmark$ $\checkmark$
++ Update
+Rsum-LLM
+MemoChat
+COMEDY
+Theanine (ours) $\checkmark$ $\checkmark$
+w/o Relation-aware Linking
+w/o Refinement $\checkmark$
+Shuffled Timeline $\checkmark$ $\checkmark$ $\checkmark$
+Table 7: Methods considered Pareto-efficient when judged based on B-4, R-L, and Bert Score reported in Table 1. $\checkmark$ = Pareto-efficient methods.
+API costs.
+The actual API costs of all settings (ours and baselines) are in Table 8.
+Agents Cost Ratio (ours $=$ 1) Cost (per episode; \$)
+All Dialogue History 0.50 0.0067
+All Memories & $\mathcal{D}$ 0.27 0.0036
++ Update 5.71 0.0771
+Memory Retrieval 0.17 0.0023
++ Update 5.63 0.0760
+Rsum-LLM 0.42 0.0057
+MemoChat 0.52 0.0076
+COMEDY 0.61 0.0082
+Theanine (ours) 1.00 0.0135
+w/o Relation-aware Linking 0.50 0.0067
+w/o Refinement 0.71 0.0096
+Shuffled Timeline 0.20 0.0027
+Table 8: API costs for Theanine and baselines.
 Appendix J Terms for Use of Artifacts
@@
 Figure 19: Examples of Timeline Refinement and Response Generation.
+Figure 20: Theanine fails to pass TeaFarm (Example 1) - Due to sudden topic change.
+Figure 21: Theanine fails to pass TeaFarm (Example 2) - Due to sub-optimal timeline utilization during RG.
 Figure 22: The prompt for the Relation-aware memory linking.